A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1900–1909

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

e [680, 319, 409, 395, 321].

&c [493, 112].

06 [315]. 07 [638].

13th [200, 322]. 14th [322]. 15th [26, 223, 633, 632, 25, 222]. 16th [165, 272, 510, 509, 675, 271]. 17th [120, 178, 177]. 1834 [292]. 1873 [667]. 1897 [548]. 18th [72, 132, 389, 388, 131]. 1900 [25, 538, 351, 463, 683, 423]. 1901 [87]. 1903 [177]. 1904 [222, 291]. 1905 [609, 331, 618, 395, 424, 430]. 1906 [681, 271, 375, 682, 322, 470]. 1907 [454, 437, 438, 439, 536]. 1908 [542]. 19th [23, 51, 88, 286, 455, 50, 87, 454].

20th [337]. 21st [211, 198]. 27th [244].

30th [265]. 31st [23, 72, 120, 165, 198, 234, 246, 244, 286, 388, 509, 632].

4th [246].

5th [234].

9

[23, 72, 120, 165, 198, 244, 286, 388, 509, 632]. 90 [317].

A.-E [546]. Abbreviated [91]. Ability [394]. Abroad [181]. Abstract [311, 291]. according [64]. Accords [521]. Account [357]. Accounting [140, 561]. Accounts [224]. Accuracy [213]. Achievements [312]. Achille [518]. Acre [341]. Acreage [139]. Act [523, 681, 355, 227, 363]. Action [60]. Acts [572, 197, 265]. Acworth [150]. Addendum [600]. Additional [136]. Additions [12, 22, 36, 46, 58, 71, 85, 98, 109, 119, 129, 142, 154, 164, 175, 185, 194, 205, 220, 230, 240, 250, 257, 269, 281, 304, 334, 360, 383, 417, 452, 483, 505, 526, 557, 593, 627, 649, 672, 688]. Address [50, 87, 306, 172, 29, 131, 177, 454, 25, 222, 271, 77]. Adelphi [23, 72, 120, 165, 198, 244, 286, 388, 509, 632]. Administration [553, 545]. Adolphe [586]. Adolphus [412]. Adrien [375]. Advancement [306, 172, 587]. Advisable [138]. Affairs [331]. Affecting [485]. Afforestation [608, 608]. Against [641]. Age [425, 655, 64]. Agraire [408]. Agricole [408, 616]. Agricultural [34, 44, 18, 83, 96, 66, 139, 127, 116, 183, 171, 162, 218, 228, 203, 255, 267, 310, 422, 460, 532, 567, 656, 679, 330, 583, 387, 472, 157, 512, 233, 224, 456]. Agriculture [339, 347, 515, 470, 285, 299]. ai [569]. Aid [314]. Albert [521, 575, 612, 409]. Album [428]. Alcoholic [15]. Alden [438]. Alexis [408]. Alfred [294, 584]. Alkoholconsumstatistik [584]. Allatamento [569]. alle [569]. Alleged [385]. Allemagne [575, 475]. Allen [624]. Amalgamated [493]. Ambitions [298]. America [587, 42, 324]. Américaine [518]. American [140, 561, 81, 519]. Among [663]. Amount [489]. Amt [476]. Amtes [620]. Analysis [277]. Anatole [614]. Andréadès [619]. Angleterre [440, 321]. Anglo [502]. Anglo-Chinese [502]. Anhang [534]. année [536]. Annual [50, 291, 177, 568, 217, 566, 331, 25, 423, 265, 201, 23, 72, 73, 121, 120, 166, 165, 199, 198, 245, 244, 287, 286, 389, 388, 510, 509, 633, 632, 24]. Annuario [393]. Annuelle [536]. Ans [659]. Août [375]. Apercu [316]. Apercus [391]. Appendix [209]. Application [375, 40]. Applications [677]. Appointed [138, 213, 207, 208, 513]. April [476]. Arbitration [572]. Area [266]. Argument [446]. Arndt [498]. Arrangements [138]. Art [352]. Arthur [324, 588, 597]. Articles [280, 303, 333, 359, 382, 416, 451, 482, 504, 525, 556, 592, 626, 648, 671, 687]. Ashley [379]. Aspects [596]. Assistance [375]. Assize [458]. Association [306, 172, 213, 421, 140, 29, 77]. Assumptions [262]. Assurent [375]. Astronomiska [313]. Attending [426]. Auftrage [476]. August [234, 246, 491]. Auslande [574]. Aussperrungen [574]. Australasia [644]. Australasian [467, 43]. Australia [572, 538]. Australian [277]. Austria [638]. Austria-Hungary. [638]. Austrian [226]. Authoritative [538]. Authorities [135]. Autriche [575, 546]. aux [575, 609]. Auxiliary [28, 41]. Available [117, 207, 208, 76]. Avebury [377]. Average [341, 95, 69]. Averages [275]. Aves [572]. Avril [375].

B [297, 409, 431, 613, 247]. Bâtie [616]. Bailey [431]. Ballard [412]. Ballou [553]. Band [637]. Bank [619, 548]. Banker [548]. Banking [404, 477, 326]. Banks [126]. Barbour [448]. Bart [33, 454, 222]. Based [214, 463, 338, 159]. Basis [117, 207, 208]. Bates [500]. Batie [616]. Baumwolle [582]. be [138, 169]. Bearbeitet [476]. Bearing [347, 232]. Beatrice [523, 355]. Beattie [352]. Beaulieu [588, 614]. Bedeutung [581]. Been [425, 312]. Before [125]. Being [666, 291, 292, 329, 352, 201]. Beitrag [620]. Beiträge [537]. Belge [435]. Belgium [212]. Bell [293]. Bellamy [439]. Bellom [642]. Bemittelten [476]. Ben [644]. Benefices [575]. Benefit [188]. Benefits [5]. Benini [602]. Bennet [33]. Berlin [89, 104, 544].

Berliner [430]. Bertillon [395]. Berufs [637]. Berufs- [637]. Berufsstatistik [637]. Besonders [570]. Besteuerung [406]. Betriebszählung [637]. Between [312]. Beverages [15]. Beveridge [611]. Bevolkerungsbewegung [434]. Bevolkerungsstatistik [479]. Bewegung [667]. Bibliography [491]. Bienvenu [609]. Bienvenu-Martin [609]. Bill [271]. Birth [655, 485, 235, 431, 273, 39, 457, 274]. Birth-Rate [485, 235, 39]. Birth-Rates [273, 457, 274]. Births [423, 308]. bis [618, 638]. Black [402, 666]. Blaschke [343]. Block [62]. Board [636, 492, 533, 606, 682, 572, 635, 426, 200, 470]. Boards [666]. Body [573]. Bogart [520]. Bohmen [327]. Bohmert [575]. Bolard [464]. Boldt [549]. Bomboy [546]. Bon [356, 439]. Book [523, 657, 658, 296, 297, 315, 316, 318, 319, 327, 343, 344, 346, 351, 355, 369, 370, 375, 376, 401, 404, 406, 425, 428, 436, 440, 445, 476, 493, 521, 573, 574, 575, 603, 604, 607, 609, 610, 612, 614, 636, 639, 640, 641, 642, 661, 665, 666, 684, 685, 291, 325, 330, 347, 348, 356, 372, 373, 393, 399, 408, 409, 412, 413, 414, 447, 463, 468, 478, 496, 498, 502, 515, 517, 541, 542, 543, 553, 578, 582, 583, 617, 467, 568, 569, 571, 602, 331, 397, 398, 427, 441, 472, 477, 547, 548, 549, 579, 581, 587, 589, 659, 667, 683, 500, 501, 516, 518, 519, 299, 300, 320, 349, 379, 400, 437, 438, 439, 492, 495, 497, 522, 533, 540, 544, 606]. Book [662, 663, 668, 643, 644, 682, 494, 618, 624, 616, 576, 615, 294, 295, 322, 323, 371, 395, 423, 431, 435, 466, 471, 292, 293, 298, 317, 321, 324, 326, 328, 329, 352, 353, 354, 374, 377, 378, 380, 402, 403, 410, 411, 442, 443, 444, 446, 448, 470, 473, 474, 475, 499, 520, 546, 554, 572, 577, 580, 588, 590, 619, 622, 645, 646, 660, 664, 357, 424, 479, 535, 405, 465, 545, 350, 570, 585, 611, 613, 623, 635, 637, 638, 669, 290, 312, 313, 314, 345, 391, 392, 394, 396, 407, 426, 429, 430, 432, 433, 434, 449, 462, 464, 469, 513, 514, 534, 536, 537, 538, 539, 550, 551, 552, 584, 586, 605, 620, 621]. Borght [658]. Borough [523]. Boroughs [489]. both [607]. Bouniatian [517, 543]. Boverat [440]. Brabrook [560]. Bradford [29]. Bray [436]. Brazilian [539]. Bread [458]. Bremen [476]. Bremischen [476]. Brief [495, 314]. Brigade [9, 55, 106, 161, 202]. Brigstocke [328]. Britain [470, 34, 83, 127, 171, 162, 218, 228, 203, 255, 267, 266, 310, 422, 460, 532, 567, 656, 679, 487, 387, 650, 124, 378]. Britian [341]. British [397, 213, 340, 421, 111, 289, 306, 172, 29, 77, 242, 251]. Buda [93]. Budget [534]. Budgets [577, 464]. Building [16, 31, 53, 613]. Bulletin [427]. Bullock [590]. Burdens [87]. Bureau [515, 400]. Bureaux [467]. Burkett [372]. Business [519]. Byron [405].

C [657, 346, 369, 370, 330, 547, 320, 663, 317, 380, 313, 433, 253]. C.B. [597, 131, 177]. Cadbury [320]. Cadoux [544]. Calcolo [552]. Calcul [605]. Calculated [91]. Calculating [76]. Calculation [74]. Calculus [314]. Calendar [76]. Cambridge [495]. Camille [435, 391, 621]. Camillo [478]. Campagnole [609]. Canada [427, 512]. Canadian [331, 288]. Capacites [344]. Capital [181, 486, 292, 403, 650, 189]. Capitales [544]. Capitalism [354]. Care [540]. Carl [666, 582]. Carnegie [587]. Carolina [373]. Caroline [496]. Carroll [598]. Case [666, 329]. Cases [265]. Castell [331]. Causes [651, 489, 294, 295, 395, 466, 298, 577, 274]. Caxton [322]. Census [214, 7, 19, 67, 160, 528, 681, 652, 493, 3, 80, 68, 145, 463, 226, 288, 400, 159, 123, 149, 81, 363, 297, 370, 427]. Census-Taking [3]. Cent [659]. Central [573, 379]. Centres [253, 261]. Century [683, 167, 274, 275, 456]. Cereal [515]. Cereals [491]. Certain [566, 565, 492, 533, 606]. Certified [197, 265]. Ch [344]. Chômage [639, 610, 642]. Change [309]. Changes [531, 212, 95, 69, 274]. Chapin [663]. Chapman [640]. Chapter [404]. Charbon [445]. Charity [357]. Charles [684, 372, 292, 326, 442, 590]. Charlier [313]. Charts [636]. Chastin [615]. Chatterton [432]. Chatterton-Hill [432]. Child [436, 400]. Children [490, 471, 357, 426, 471]. Chinese [553, 502]. Chiozza [353]. Chiozza-Money [353]. Christabel [619]. Christian [604]. Cinquantenaire [680]. Circulaires [375]. Circulation [654]. Cities [60]. Citta [393]. City [357, 320, 663, 464]. Civil [92, 535, 536]. Clarence [372]. Clark [500]. Class [492, 533, 606]. Classe [399]. Classes [60]. Classification [309, 466]. Clay [588]. Clementina [666, 402]. Clothes [666]. Clothing [682]. Co [404, 467, 404]. Co-Operative [404, 404]. Co-Ordination [467]. Coal [340, 189, 350, 167]. Coal-Capital [189]. Coast [608]. Coefficient [639]. Collection [138, 426]. Collectivism [588]. Colleges [224]. Collings [413]. Colonel [598]. Colonies [43]. Columbia [586]. Combination [398]. Comfort [601]. Command [607]. Commencements [321]. Commentaire [375, 521, 609]. Commercants [643]. Commerce [316, 502, 643, 17, 513]. Commercial [8, 54, 105, 151, 191, 238]. Commission [608, 607, 665, 472, 540, 401]. Commissioners [542]. Committee [138, 117, 213, 207, 208, 528, 652, 573, 541, 542, 471, 572, 513, 227, 472]. Commodities [6, 52, 102, 147, 190, 237, 278, 366, 488, 599]. Common [414]. Commonwealth [467, 538]. Company [396, 448, 554]. Comparability [213]. Comparative [324]. Comparatives [544]. Comparee [316]. Comparison [674, 676]. Comparisons [339, 341]. Compensation [227, 197, 265]. Complete [292]. Completed [265]. Compte [296]. Conant [326]. Concerning [350]. Conciliation [572]. Concluded [53]. Condition [357, 426]. Conditions [595, 431, 471, 251, 578, 636]. Condizioni [399, 569]. Conference [322]. Conferences [391]. Congress [419, 461, 491, 259, 93]. Connected [243]. Connection [425]. Conrad [668]. Consideration [468]. Considerations [418]. Considered [112]. Consignment [291]. Consols [548]. Constance [665]. Constants [511, 529, 563, 600]. Constructing [136]. Construction [657, 40]. Constructive [589]. Consumption [117, 209, 207, 208, 630, 15, 468, 210]. Containing [538]. Contd [31, 529, 563]. Contenant [375]. Continent [122]. Continuous [28, 41]. Contrat [614]. Contre [610, 642]. Contribution [639, 252]. Control [540]. Copenhagen [491]. Corn [89]. Cornélissen [604]. Corporations [523, 355]. Corrected [273, 538]. Correction [104]. Correlation [79, 260, 362, 677, 369, 462]. Correspondence [301, 480, 368]. Corresponding [76]. Cost [595, 464]. Cotton [64, 372, 373, 284, 372]. Council [23, 120, 165, 198, 244, 286, 388, 509, 632]. Countries [291, 277, 273]. Country [578]. County [355]. Cours [536]. Credit [516, 404]. Criminal [307]. Crise [408]. Crises [374]. Crisi [478]. Critical [232, 589]. Crops [341, 362]. Crozier [352]. Cultivation [372]. Currency [477]. Curves [28, 41, 369]. Custody [138].

D [614, 598, 398, 328, 261]. D.C.L [50, 87]. d’Administration [375]. Dairy [210]. Dangers [441]. dans [349, 435]. Danvers [134]. d’après [535, 536]. Darien [448]. Data [232]. d’Avenel [659]. Davenport [589]. David [641, 624, 469]. Dawson [318, 427, 623]. Day [588]. Dealings [284]. Dearle [613]. Death [431, 466]. Deaths [214, 159, 423]. Decade [168, 182, 100].

December [23, 72, 120, 165, 198, 244, 286, 388, 509, 632]. Decennium [214, 91]. Deces [395]. Decline [387, 273]. Deduced [265]. Defects [471]. Defence [27]. Degrees [312]. Delivered [50, 87, 131, 177, 25, 222, 271, 454]. Della [393, 399]. Demografiche [552]. Demographie [435]. Denmark [285]. Department [515, 572, 513]. Depopulation [347, 294, 295]. Depuis [659]. Deputation [200, 476]. Derived [463]. Designed [314]. Destination [291]. Detailed [291]. d’Etat [535, 536]. Deterioration [396]. Determining [462]. Deutsche [497, 429]. Deutschen [327, 581, 637, 669, 327]. Deutschland [406, 434]. Devant [408]. Devas [442]. Development [512, 285, 271, 622]. Dewsnup [437]. Diathesis [469]. Dietzel [328]. Different [76]. Differential [459]. Diplomacy [502]. Directed [251]. Direction [167]. Directs [645]. Disclosed [90]. Discussion [560, 421]. Disease [431, 466]. Diseases [571]. Dissolution [297]. Distress [607, 665]. Distribution [346, 225, 589, 215, 235, 38]. Districts [689, 308, 187]. Divorce [431]. Dockyards [261]. Domesday [412]. Domestic [531]. Donation [342]. Dondlinger [617]. Données [391]. Door [489]. Douglas [519]. Dr [668]. Dr. [562, 390]. Draper [396]. Dressmaking [666]. Dupont [610]. During [31, 103, 236, 247, 253, 261, 233, 167, 456, 265, 16, 53, 275, 243, 157, 531, 274]. Duties [328, 380]. Duty [338, 486].

Each [565, 91]. Early [554]. Earned [664]. Earner [464]. Earners [427]. Earnings [682, 635, 682, 635]. Eberstadt [661]. ecc [399]. Economic [229, 239, 249, 256, 268, 279, 280, 303, 333, 359, 382, 416, 451, 482, 504, 525, 556, 592, 626, 648, 671, 687, 306, 500, 77, 298, 520, 622, 596, 677, 172, 29, 317, 646]. Economical [11, 21, 35, 45, 57, 70, 84, 97, 108, 118, 128, 141, 153, 163, 174, 184, 193, 204, 219]. Economiche [569, 478]. Economics [262, 314, 586, 376, 330, 590]. Economie [444]. Economique [443]. Economiques [294]. Economist [82, 61]. Economy [292, 352, 293, 442]. Edgar [312, 394]. Edinburgh [357]. Edmond [616]. Edouard [609]. Education [413]. Edw [320]. Edward [560, 494]. Edwin [299]. Effect [89, 197, 251]. Effects [325]. Efficiency [324, 318, 624]. Eglantyne [495]. Eidgenössischen [620]. Eighth [121, 120, 93]. Elderton [369, 551]. Electoral [336]. Élémentaires [605]. Elements [590, 290]. Elis [497]. Eltzbacher [298]. Empire [145, 289, 27, 169, 533, 553]. Employed [493]. Employment [629, 370]. Enclosure [414]. Ended [23, 72, 120, 165, 198, 244, 286, 388, 509, 632]. Ending [23, 72, 120, 165, 198, 286, 388, 509, 632, 265, 244]. Energique [344]. Engelbrecht [618]. Engineering [247, 253, 236, 247, 253, 261, 275]. England [406, 447, 568, 437, 423, 324, 622, 689, 157, 91, 159, 283, 206, 456, 308, 38, 90, 187, 457, 274, 543, 619]. Englischen [406]. English [297, 414, 583, 133, 10, 56, 107, 152, 192, 339, 16, 158, 248, 523, 355]. Enquête [616]. Enquiry [492, 533]. Enterprise [519, 25]. Entwicklung [658, 581]. Enumerations [214, 159]. Envisagé [612]. Epstein [554]. Ernest [520, 572]. Ernst [343, 407]. Erosion [608]. Errata [689]. Errichtung [620]. Error [113, 305]. Errors [511, 529, 563, 600]. Especial [197]. Especially [314]. Essai [321]. Essais [621]. Essay [674]. Essays [545]. Essentials [500]. Estate [338, 486]. Estimated [341]. Estimates [338, 233, 135]. Estimating [117, 207, 208, 654, 276, 486, 63]. Ethel [551]. Etude [435, 639]. Études [544]. Eugen [406]. Eugenics [392]. Europe [515, 674]. European [515]. Evidence [573].

Evolution [522, 616, 354, 623, 352]. Examination [92, 651, 489, 232, 471]. Exchange [89, 104]. Exhaustion [350]. Existing [138, 447]. Expenditure [50, 27, 188, 206, 5]. Experience [197]. Exports [309, 351, 291, 424, 189, 468]. Expropriations [662]. Exterieur [316]. Extracts [213].

F [657, 641, 325, 345]. F.I.A. [634]. F.R.S [50, 87]. F.R.S. [33]. Fachleute [407]. Factors [485]. Factory [112]. Facts [90]. Faculty [586]. Fairlie [545]. Falk [370]. Falkenburg [570]. Fall [548]. Families [663, 312]. Farm [42]. Fascicule [344]. Faure [290]. February [51, 50]. Federal [144]. Feeble [540]. Feeble-Minded [540]. Felix [444]. Femme [349]. Fernand [290]. Ferrand [356]. Fertility [273]. Field [373]. Fields [414]. Fifth [633, 632, 568, 566, 423]. Fifty [157, 531]. Finance [144, 25]. Financial [23, 72, 120, 165, 198, 244, 286, 388, 509, 632, 587, 101]. Financier [612]. Financiers [440]. Fire [9, 55, 106, 161, 202]. Fires [9, 55, 106, 161, 202]. First [67, 245, 244, 297, 396, 469]. Fiscale [546]. Fisher [501, 403, 314]. Fisheries [470]. Fishery [386]. Five [197]. Flanders [435]. Flux [350]. Fondate [552]. Food [508]. Foreign [213, 340, 339, 311, 30, 277, 20, 17, 167]. Forestry [596]. Formulae [4]. Fortune [616]. Fortunes [576]. Forty [367]. Foster [426]. Found [471]. Foundation [587, 392]. Foundations [293]. Fourteen [357]. Fourth [510, 509, 201]. Foville [264, 474]. Fox [597, 170]. Foxwell [619]. Frage [584, 620]. Fran [313]. France [576, 473, 577, 535, 536, 30, 606, 508, 445, 496, 428]. Francis [312, 392]. Francois [445]. Frank [685, 586]. Franz [603]. Frau [497]. Frederick [634]. Free [329]. French [181]. Frequence [395]. Frequency [365, 511, 529, 563, 600, 369].

Frequency-Constants [511, 529, 563, 600]. Frequency-Curves [369]. Friday [200]. Friedrich [371]. Friendly [243]. Fulfilling [113]. Fürth [534]. Further [138, 217, 462]. Future [263]. Futures [284].

G [657, 342, 370, 659, 320, 494, 390, 466, 353, 402, 538, 551]. G.C.M.G. [271]. G.C.S.I [25]. Gainsborough [401]. Galton [312, 392]. Gardiner [402]. Gaspard [475]. Gaston [544]. Gauthier [546]. General [23, 72, 73, 121, 120, 166, 165, 199, 198, 245, 244, 287, 286, 389, 388, 510, 509, 633, 632, 24, 247, 275, 568, 516, 77, 423, 217, 566, 201]. Generale [428, 535]. Generales [374]. Generali [399]. Generalised [305]. Generaux [391]. Genitori [569]. Geographical [225]. Geographische [618]. George [414, 468, 323, 432, 667]. German [318, 533, 126]. Germany [587, 298, 324, 20, 508, 401, 623]. Geschichte [543]. Gesetzgebung [620]. Getreidepreise [618]. Gewerbe [638]. Ghent [64]. Gibson [548, 548]. Giornate [399]. Giving [493]. Glasgow [77, 426]. Gnauck [497]. Gnauck-Kuhne [497]. Gold [75, 405, 547]. Gonnard [349]. Gonner [317]. Goschen [364]. Gottlieb [498]. Government [523, 200, 355, 636, 144, 379, 539]. Grain [515]. Grande [321]. Grandes [544]. Graphique [428]. Gray [297, 550]. Great [305, 378, 470, 34, 83, 127, 171, 162, 218, 228, 203, 255, 267, 266, 310, 341, 422, 460, 532, 567, 656, 679, 487, 387, 650, 124]. Greenback [547]. Grève [614]. Grigaut [444]. Groups [113]. Growth [674, 50, 683, 180, 167]. Grundriss [668]. Guarantee [542]. Gustav [479]. Guyot [643, 443].

H [319, 370, 640, 542, 398, 548, 589, 349, 439, 644, 618, 624, 466, 328, 619, 622, 465, 570, 611, 514, 538, 586]. Habitations [439].

Halbwachs [662]. Half [167, 274]. Hall [322, 325]. Halle [407]. Hallsworth [640]. Ham [494]. Hamilton [372]. Handbuch [429, 661, 371]. Handel [582]. Handelskrisen [543]. Harbutt [318, 623]. Hardy [657]. Hasbach [583]. Hauser [319]. Hawkins [586]. Hayward [438]. Health [636, 232]. Heinrich [327]. Held [23, 72, 120, 198, 286, 389, 388, 510, 509, 29, 77, 322, 165, 244, 633, 632, 172]. Henderson [684]. Hendriks [634]. Henriette [534]. Henry [344, 376, 404, 330, 397, 477, 522, 576]. Her [298]. Herbert [392]. Heredity [345, 432]. Heresies [646]. Heron [469]. Herring [386]. High [651]. Hill [317, 432]. Historical [39]. Historique [521]. History [8, 54, 105, 151, 191, 238, 297, 583, 448, 520, 554, 619, 622, 586, 446, 300]. Hobson [354, 664]. Holding [418]. Holdings [447]. Holland [373]. Holmes [468]. Holt [405]. Home [471, 572]. Homes [357]. Hon [50, 87, 454, 25, 271]. Honourable [312]. Hop [541]. Hopkins [331]. Hosea [553]. Hospital [595]. Hours [493, 682, 635]. Houses [607]. Housing [437, 492, 533, 606, 60, 438]. Howarth [494]. Hugo [378]. Hull [477]. Human [273]. Hundred [16, 31, 53, 103, 236, 247, 253, 261, 357]. Hungary [578]. Hungary. [638]. Hygienic [251].

Ideas [77]. Igier [296]. II [291, 682, 313, 396, 566]. Illustrated [39]. im [476, 574, 638, 534]. Immigration [431, 325]. Impôt [546]. Impôts [645]. Imperial [411, 87]. Import [380]. Importance [77]. Imports [309, 291, 277, 424, 351]. Improvement [678, 527, 232, 251]. In- [574]. In-Door [489]. Inama [562]. Inama-Sternegg [562]. Inaugural [87, 131]. Incapacity [265]. Inchiesta [569]. Income [111, 403, 664].

Increase [630, 385]. Incurables [375, 609]. Indebtedness [206]. Index [275, 173, 312]. India [111, 148, 651, 225, 125, 411]. Indian [410, 68, 507, 49]. Indien [618]. Indirects [645]. Industrial [685, 373, 519, 320, 492, 533, 606, 494, 324, 410, 572, 664, 251, 684, 398, 324]. Industrie [321]. Industrielle [321]. Industrielles [473]. Industry [64, 122, 644, 402, 397, 611, 541]. Infant [564, 323]. Infantile [569, 322, 435]. Infection [551]. Infinitesimal [314]. Infirmes [375, 609]. Influence [276, 48]. Information [1, 37, 47, 86, 99, 130, 143, 176, 186, 221, 231, 258, 270, 335, 361, 453, 484, 558, 594, 673, 463]. Ingenbleek [645]. Inheritance [394, 469]. Injury [265]. Inquest [412]. Inquire [138, 117, 213, 207, 208, 513]. Inquiries [655]. Inquiry [682, 664, 350, 456, 635, 274, 606]. Insane [469]. Insanity [385, 469]. Instability [125]. Institut [344, 391, 680]. Institute [234, 246, 419, 491, 653, 216, 254, 93]. Institutions [243]. Instituts [344]. Instructions [375]. Instruments [516]. Insurance [684, 685, 641]. Insurances [559]. Intellectuelle [344]. Inter [346]. Inter-Temporary [346]. Intercensal [76, 63]. Interest [501]. International [213, 461, 674, 680, 676, 2, 234, 246, 419, 491, 653, 216, 254, 259, 93]. Internationale [535]. Internationaux [521]. Introduction [330, 314, 248, 376]. Investigation [640, 666, 137]. Investments [548, 181, 650]. Ireland [31]. Irish [424]. Irving [501, 403, 314]. Isles [242]. Issued [539]. Italiane [393]. Italy [156]. IX [103].

J [370, 640, 347, 502, 331, 579, 589, 668, 615, 446, 664, 539]. Jacob [577]. Jacquart [435, 391, 621]. Jacques [395]. Jahr [407, 430]. Jahr-und [407]. Jahresperioden [638].

Jahrhundertwende [497]. James [346, 448]. January [271, 272]. Jean [374]. Jebb [447, 495]. Jephson [522]. Jesse [413, 590]. Jevons [350]. Jews [252]. John [33, 500, 292, 317, 352, 354, 580, 545, 550]. Jones [665]. Joseph [577]. Jours [577]. Jouve [612]. Juillet [375, 609]. Juland [134]. Jules [348, 356, 645]. Julius [578]. July [234, 246]. June [322, 23, 72, 120, 165, 211, 198, 244, 286, 389, 388, 510, 509, 633, 632, 265]. Juni [637]. Jur [549]. Justin [348]. Juvenile [14].

Karl [396, 462, 551]. Kemmerer [516]. Kindersterblichkeit [570]. King [579]. Kingdom [291, 540, 682, 418, 44, 96, 117, 209, 207, 208, 16, 31, 53, 103, 236, 247, 253, 261, 275, 144, 492, 338, 277, 631, 273, 233, 252, 508, 60, 635, 493]. Kinships [312]. Kirkman [297]. Klassen [476]. Knibbs [466, 538]. Knoop [519]. Kolonien [406]. Kommunal [406]. Kommunal-Besteuerung [406]. Kraft [585]. Kubatz [584]. Kuhlmann [582]. Kuhne [497]. Kulturgeschichtliche [585].

L [447, 353, 313]. Labor [400, 513]. Labour [311, 401, 493, 573, 682, 635]. Labourer [583]. Labourers [387]. Lancashire [640]. Land [456, 348, 413]. Lands [650]. Language [248]. Large [60]. L’Argentine [409]. L’Assistance [609]. L’Assurance [610, 642]. Last [16, 31, 53, 103, 236, 247, 253, 261, 157, 233, 167, 63]. Late [364]. Laurent [514]. Lauterbach [406]. Lavergne [576]. Lavoro [399]. Law [113, 305, 459, 305, 586]. Lawes [33]. Laws [607, 665]. Layton [657, 369, 657, 369]. Lazard [639]. Leading [651, 588]. Lebenshaltung [534]. Lecons [605]. Lecture [392]. Led [489]. Legislation [112, 271]. Legislative [571]. Lehre [343]. Length [565]. Leroy [588, 614]. Leroy-Beaulieu [614]. Lescure [374].

Lesebuch [407]. Leslie [426]. Lethbridge [411]. Levant [554]. Levasseur [316, 473, 660]. Level [651]. Lewandowski [409]. Lewis [685]. L’Habitation [356]. Library [12, 22, 36, 46, 58, 71, 85, 98, 109, 119, 129, 142, 154, 164, 175, 185, 194, 205, 220, 230, 240, 250, 257, 269, 281, 304, 334, 360, 383, 417, 452, 483, 505, 526, 557, 593, 627, 649, 672, 688]. Lichte [534]. Life [214, 507, 564, 320, 40, 91, 114, 136, 159, 629, 324, 401]. Life-Table [214, 507, 159]. Life-Tables [40, 91, 114, 136]. Limitations [3]. L’Impôt [546]. L’Impot [475]. L’Industrie [349, 518]. Lines [352]. List [12, 22, 36, 46, 58, 71, 85, 98, 109, 119, 129, 142, 154, 164, 175, 185, 194, 205, 220, 230, 240, 250, 257, 269, 281, 304, 334, 360, 383, 417, 452, 483, 505, 526, 557, 593, 627, 649, 672, 688]. Literature [56, 107, 152, 192]. Literture [10]. Live [139]. Living [663, 464]. LL.D. [50, 87]. Local [523, 138, 87, 355, 636, 206, 271, 200, 379]. Loch [370]. Loi [375, 609]. London [573, 666, 613, 9, 23, 55, 72, 106, 120, 161, 165, 202, 198, 234, 246, 244, 286, 388, 509, 632, 53, 595, 168, 182, 196, 146, 60, 123, 522, 315]. Londres [544]. Lord [50, 87, 377]. Lords [542]. Losses [233]. Louise [464]. Louvain [391]. L’Ouvriere [496]. Lucien [356, 295]. Ludlow [520]. Luigi [552]. Lunds [313]. Lycée [615].

M [370, 609, 614, 640, 356, 542, 579, 320, 390, 294, 295, 444, 554, 580, 551]. M. [62, 264]. M.P. [454, 25, 222]. Macgregor [398]. Machine [354]. Machtentfaltung [585]. Mackenzie [426]. Macrosty [397]. Made [135, 640]. Maize [42]. Majesty [607]. Major [131, 177]. Making [666]. Manor [523]. Mantoux [321]. Manufacture [372]. Manufactures [103]. Marcé [546]. March [295]. Marche [439, 356]. Marital [551]. Market [284, 2].

Marketing [372]. Marriage [79, 431, 457, 274]. Marriage- [274]. Marriage-Rate [79]. Marriage-Rates [457]. Marriages [423]. Married [235]. Martel [444]. Martin [609]. Martinez [409]. Mary [317]. Massenkonsums [534]. Masszahlen [343]. Mathematical [4, 314]. Mathématique [514]. Mathematische [343]. Matheson [320]. Maurice [62, 296, 642, 409, 662]. Max [639]. Maximilian [574]. May [25, 138, 26, 200]. McCarthy [348]. Mean [76, 63]. Measures [138, 179]. Meat [117, 209, 207, 208, 631, 135, 210, 468]. Mecum [548]. Meddelanden [313]. Median [676]. Medical [571]. Medizinischen [371]. Meeker [300]. Meeting [211, 421, 23, 72, 73, 121, 120, 166, 165, 199, 198, 245, 244, 287, 286, 389, 388, 510, 509, 633, 632, 24]. Meline [348]. Memoir [550]. Memoires [344]. Memoirs [396]. Memoranda [636]. Memorandum [291, 135, 149, 158, 457, 425]. Men [493, 663]. Mendelism [433]. Mental [345]. Mentor [517, 543]. Merchandise [277]. Meredith [619, 622]. Mesure [344]. Method [654, 91, 136, 486, 63, 677]. Methodologische [603]. Methods [421, 561, 467, 232, 113, 76, 462]. Métin [521]. Metodologica [602]. Metrical [179]. Metropolitan [489, 9, 55, 106, 161, 202]. Meyer [666, 574, 378]. Michel [616]. Milano [399, 569]. Milhaud [496]. Milk [117, 209, 207, 208, 135]. Mill [373]. Minded [540]. Minder [476]. Mines [445, 350]. Minimum [402]. Ministérielles [375]. Minutes [573]. Mitchell [547]. Mittelbürgerliches [534]. Mittelwerte [603]. Mitwirkung [407]. Mixter [292]. Modern [421, 352, 354, 623, 312, 431, 298]. Moderne [321]. Modi [569]. Moitie [395]. Mombert [434]. Monasteries [297]. Money [654, 326, 2, 353, 516]. Monnaie [474]. Monopoly [48]. Montessus [605].

Moral [345]. Morale [621]. Morgan [644]. Morison [410]. Morse [553]. Mortalita [569]. Mortalite [435]. Mortality [65, 420, 490, 322, 323, 657]. Mortgage [404]. Most [76]. Mostly [338]. Mouvement [535, 536]. Movement [397]. Movements [30]. Mr. [342, 170, 150]. Municipal [523, 50, 355, 440, 441, 25, 377, 545, 60, 378, 499].

N [620]. Nathan [646]. Nathaniel [646]. Nati [569]. Nation [350]. National [50, 318, 630, 74, 542, 322, 377, 396, 322]. Nations [446]. Nature [403]. Near [312]. Nearly [76]. nebst [534]. Need [685]. nel [569]. Newman [323]. Neymarck [294]. Nicholson [347]. Niederlanden [570]. Nineteenth [275, 456]. Ninety [38]. Ninth [166, 165]. Non [247, 253, 616]. Non-Trade [247, 253]. Nordamerika [406]. Nordamerikanischer [582]. Normal [113]. Norman [317, 613]. North [373, 565, 512]. North-West [512]. nos [577]. Notamment [395]. Note [655, 140, 217, 420, 566, 565, 276, 215, 235, 490, 201, 81]. Notes [11, 21, 35, 45, 82, 57, 70, 84, 97, 108, 118, 128, 141, 153, 163, 174, 184, 193, 204, 219, 229, 239, 249, 256, 268, 279, 344, 80, 170, 114, 136, 150, 61, 680]. Noteworthy [312]. November [87, 177, 454, 222, 88, 132, 178, 223, 337, 455, 675, 131]. Nr [313]. Number [565, 387]. Numbers [14, 275, 76, 173, 305].

O [298, 622]. Obligatoire [609]. Observations [421, 260, 210]. Observatorium [313]. Occasion [680]. Occupation [571]. Occupations [226, 492, 533, 606, 215, 251, 158, 427]. Occupiers [530]. Occupying [413]. Oekonomie [668]. Oesterreichischen [465]. Offenders [14]. Office [243]. Officers [140]. Official [678, 527, 538]. Officiels [521]. Okonomische [585]. Olanesco [390]. Old [425, 655, 425]. Oliver [571]. Operaia [399]. Operation [499]. Operative [404, 404]. Operatives [112]. Operees [296]. Opinion [94]. Ordination [467]. Organisation [410, 357, 519]. Organized [60]. Osborne [328]. Other [302, 332, 358, 381, 415, 450, 481, 503, 524, 555, 591, 625, 647, 670, 686, 243, 113, 548, 273, 650, 252, 60]. Our [189, 167, 666, 350]. Outlines [622]. outre [375]. Ouvrieres [473, 612]. Ouvriers [445, 521, 319]. Ownership [413, 378]. Oxford [458, 224].

P [617, 441, 438, 539]. P.C [87]. Palin [369]. Paper [560, 170, 150]. Papers [449]. Paris [544, 662, 395, 653, 680]. Parish [355]. Parliament [607]. Part [565, 16, 31, 53, 103, 236, 247, 253, 261, 275, 566]. Participation [575]. Passe [319]. Past [274]. Paterson [448]. Patron [296]. Patronage [539]. Paul [610, 667, 576, 321, 588, 434]. Pauperism [168, 182, 489, 283]. Pearson [551, 396, 462]. Peasant [413]. Peasantry [414]. Pendant [395]. Pendulum [336]. Pensions [425, 655]. People [578]. Peradventures [507]. Percy [379]. Period [76, 63, 38, 538]. Periode [395]. Periodical [13, 59, 110, 155, 195, 241, 282, 384, 506, 628]. Periodicals [280, 303, 333, 359, 382, 416, 451, 482, 504, 525, 556, 592, 626, 648, 671, 687]. Periodiques [374]. Perozzo [552]. Persia [82, 61]. Persons [312]. Pesth [93]. Ph [570]. Phase [320]. Philanthropy [297]. Philosophie [391]. Physical [471, 357, 426, 471]. Physicians [471]. Pictures [370]. Pigmentation [550]. Pigou [380]. Plaice [565]. Point [612]. Points [571]. Policy [298]. Political [292, 293, 298, 352, 586, 442]. Politique [444]. Politischen [668]. Poor [607, 665]. Pope [551]. Population [674, 225, 289, 683, 76, 283, 535, 48, 39, 38, 90, 536]. Populations [63]. Porter [441]. Portions [123]. Ports [424]. Position [418]. Pour [296]. Poverty [353]. Practical [654, 40, 477]. Practice [404, 81]. Pratt [299]. Prefatory [291]. Preference [411]. Preferential [380]. Preliminary [425, 573, 90, 341, 681]. Prepared [425, 636]. Prescott [325]. Present [651]. Presented [23, 72, 120, 165, 198, 244, 286, 388, 509, 632, 607, 491]. President [364, 200, 50, 87, 131, 177, 25]. Presidential [454, 222, 271]. Prevailing [492]. Previous [655]. Prévoyance [521]. Price [30]. Prices [491, 148, 651, 351, 547, 492, 533, 606, 42, 89, 125, 6, 52, 102, 173, 147, 190, 237, 278, 366, 488, 599, 124, 95, 137, 516]. Primary [158]. Principal [341, 492, 533, 606, 158]. Principales [395]. Principii [602]. Principles [404, 292, 326]. Prinzing [371]. Private [499, 60, 271]. Privée [616]. Prives [576]. Prix [662]. Probabilités [605]. Probability [342, 313, 392]. Probable [511, 529, 563, 600, 350, 274]. Problem [437, 611, 537, 323]. Problème [612]. Problems [372, 477, 298, 494, 613]. Proceedings [26, 51, 88, 73, 132, 121, 178, 166, 200, 211, 223, 199, 245, 272, 337, 287, 455, 389, 510, 675, 633, 24, 93, 322, 572]. Produce [341, 346, 89, 104]. Producteurs [615]. Production [117, 209, 207, 208, 681, 515, 75, 470, 135, 210, 363, 681, 363, 354]. Products [210]. Prof [668]. Profession [639]. Professionel [639]. Professionnelles [295]. Professor [587, 32]. Profits [347]. Progress [243, 43, 367, 350, 38]. Progression [577]. Proposed [7, 19]. Proprietary [413]. Propriété [616]. Prosperity [48, 405]. Protected [112]. Protection [329]. Protective [380]. Province [410]. Provinces [512]. Prusse [546]. Public [636, 232, 331, 499, 426, 586]. Publications [302, 332, 358, 381, 415, 450, 481, 503, 524, 555, 591, 625, 647, 670, 686]. Publicly [312]. Publique [375]. Publishing [232]. Pulmonary [396, 551]. Punnett [433]. Purpose [138].

Quarterly [12, 22, 36, 46, 58, 71, 85, 98, 109, 119, 129, 142, 154, 164, 175, 185, 194, 205, 220, 230, 240, 250, 257, 269, 281, 304, 334, 360, 383, 417, 452, 483, 505, 526, 557, 593, 627, 649, 672, 688]. Question [425, 350]. Questions [588, 495, 473]. Quetelet [586]. qui [375]. Quinquennial [7].

R [658, 437, 570, 433, 605]. Rae [292]. Railway [133, 493, 261, 140, 561]. Railways [493, 487]. Rapport [294, 295]. Rapporto [569]. Rate [485, 548, 501, 79, 235, 39]. Rates [236, 180, 492, 533, 606, 273, 457, 274]. Ratios [308]. Rauchberg [327, 465]. Raymond [440]. Re [329]. Re-Statement [329]. Reading [314]. Real [212, 601]. Realisable [338]. Received [681]. Reconstruction [352]. Recorded [214, 159, 312]. Recording [232]. Records [138]. Reddito [399]. Reference [431, 224, 197, 112]. Reform [413]. Réforme [546]. Reginald [436]. Registrar [568, 423, 217, 566, 201]. Registrar-General [568, 423, 217, 566, 201]. Registration [655, 308]. Registres [535, 536]. Registry [243]. Reglements [375]. Regulations [251]. Reich [429]. Reichesberg [620]. Reichs [637, 669]. Reichsfinanzen [658]. Relating [418, 636]. Relation [516, 235]. Relations [347]. Relative [276]. Relief [607, 665, 595]. Remèdes [577]. René [546]. Rent [456]. Rents [347, 492, 533, 606]. Reorganisation [513]. Report [18, 23, 66, 116, 138, 120, 117, 162, 165, 203, 213, 207, 208, 198, 244, 286, 388, 528, 509, 608, 652, 632, 493, 573, 607, 93, 568, 217, 566, 472, 227, 540, 682, 322, 423, 470, 572, 90, 513, 201, 541, 542, 492, 533, 606, 572, 357, 424, 426]. Reports [209, 624]. Representation [4, 365]. Representing [113]. Reprint [292]. Republique [473]. Research [396]. Researches [313]. Resources [74].

Resultats [316, 440]. Results [67, 681, 493, 640, 666, 275, 595, 288, 60]. Retail [492, 533, 606, 137]. Retaliatory [328]. Retenues [296]. Retraites [612]. Retrospect [101]. Return [348, 329]. Returns [34, 44, 13, 18, 83, 96, 59, 66, 139, 127, 110, 116, 183, 171, 162, 155, 218, 228, 203, 195, 255, 267, 241, 310, 282, 422, 460, 384, 532, 567, 506, 656, 679, 681, 628, 338, 265]. Revenu [475, 546, 546, 645]. Revenue [579]. Revenues [224]. Revenus [576]. Review [523, 657, 658, 8, 54, 105, 151, 191, 238, 111, 296, 297, 315, 316, 318, 319, 327, 343, 344, 346, 351, 355, 369, 370, 375, 376, 401, 404, 406, 425, 428, 436, 440, 445, 476, 493, 521, 573, 574, 575, 603, 604, 607, 609, 610, 612, 614, 636, 639, 640, 641, 642, 661, 665, 666, 684, 685, 291, 325, 330, 347, 348, 356, 372, 373, 393, 399, 408, 409, 412, 413, 414, 447, 463, 468, 478, 496, 498, 502, 515, 517, 541, 542, 543, 553, 578, 582, 583, 617, 467, 568, 569, 571, 49, 602, 331, 397, 398, 427, 441, 472, 477, 547, 548, 549, 579, 581, 587, 589, 659, 667, 683, 500, 501, 516, 518, 519, 299, 300, 320, 349, 379]. Review [400, 437, 438, 439, 492, 495, 497, 522, 533, 540, 544, 606, 662, 663, 668, 643, 644, 682, 494, 618, 624, 616, 576, 615, 294, 295, 322, 323, 371, 395, 423, 431, 435, 466, 471, 292, 293, 298, 317, 321, 324, 326, 328, 329, 352, 353, 354, 374, 377, 378, 380, 402, 403, 410, 411, 442, 443, 444, 446, 448, 470, 473, 474, 475, 499, 520, 546, 554, 572, 577, 580, 588, 590, 619, 622, 645, 646, 660, 664, 357, 424, 479, 535, 405, 465, 545, 350, 570, 585, 611, 613, 623, 635, 637, 638, 669, 290, 312, 313, 314, 345, 391, 392, 394, 396, 407, 426, 429, 430, 432, 433, 434, 449, 462, 464, 469, 513, 514, 534, 536, 537, 538, 539, 550, 551, 552]. Review [584, 586, 605, 620, 621]. Revolution [523, 355, 321]. Reyer [585]. Ricerche [552]. Richard [537]. Riches [659, 353]. Richesse [576]. Richmond [684]. Riding [103]. Right [50, 87, 454, 25, 271].

Rise [271]. Risks [542]. Risultati [569]. Robert [441, 663, 378]. Robertson [293, 580]. Rodolfo [602]. Rogers [376]. Rooms [23, 72, 120, 165, 198, 244, 286, 388, 509, 632]. Roper [411]. Roumania [367]. Royal [50, 87, 131, 177, 25, 300, 540, 23, 72, 120, 165, 198, 244, 286, 388, 491, 509, 608, 632, 454, 222, 607, 665]. Royalty [345]. Rud [661]. Rupee [148, 651]. Rural [347, 413]. Russia [515, 48]. Russie [408].

S [370, 640, 347, 427, 624, 442, 619]. Sachet [375]. Salaire [445, 604]. Salaires [296, 614, 660]. Salari [399]. Salariat [660]. Salarié [604]. Saltet [570]. Samples [565]. Samuel [448]. Sanitary [522]. Sargent [502]. Schedules [463]. Schemes [197, 265]. Schloss [641]. Schnapper [498]. Schnapper-Arndt [498]. School [624, 471, 357, 426, 624]. Schools [426]. Schuchart [581]. Schuster [312, 394]. Schwedens [479]. Science [306, 172, 29, 77, 352, 443, 312, 391, 586, 306, 172]. Scotland [214, 655, 31, 550]. Sea [565]. Seager [376]. Seasons [242]. Second [287, 286, 573, 551, 207, 608]. Seconde [395]. Secretary [572, 513]. Section [306, 77, 172, 29]. seit [667]. Select [541, 572]. Selection [432]. Sept [659]. Serie [313]. Series [365, 91]. Servants [493, 531]. Service [92]. Ses [577, 440]. Session [234, 246, 680, 50, 87, 131, 177, 454, 25]. Sessional [23, 72, 120, 165, 198, 244, 286, 388, 509, 632]. Settlement [512]. Seventh [72, 73]. Seventieth [199, 198]. Seventy [245, 244, 287, 286, 389, 388, 510, 509, 633, 632]. Seventy-Fifth [633, 632]. Seventy-First [245, 244]. Seventy-Fourth [510, 509]. Seventy-Second [287, 286]. Seventy-Third [389, 388]. Sex [64, 308]. Sex-Ratios [308]. Shadwell [324]. Shann [320]. Shipbuilding [236, 247, 253, 261, 275]. Shipping [542, 300, 78]. Short [136]. Should [169]. Showing [341, 291]. Shown [251, 273]. Sidgwick [32]. Sidney [523, 355]. Sieben [638]. Siecle [395, 321, 577, 409]. Siecles [319]. Siegfried [356]. Silberpreises [667]. Simiand [445]. Simon [328]. since [548, 5, 601]. Sir [454, 25, 33, 134, 560, 222]. Sixth [23, 24]. Sixty [23, 72, 73, 121, 120, 166, 165, 24, 568, 566, 423, 201]. Sixty-Eighth [121, 120]. Sixty-Fifth [568, 566, 423]. Sixty-Fourth [201]. Sixty-Ninth [166, 165]. Sixty-Seventh [72, 73]. Sixty-Sixth [23, 24]. Sketch [578]. Skimmings [259]. Slater [414]. Small [418, 447]. Smaller [689, 187]. Smart [329]. Smith [346]. Snedden [624]. Social [636, 685, 559, 530, 571, 495, 494, 323, 431, 588, 622, 677]. Sociale [391]. Socialisme [440]. Société [680, 680]. Societies [243, 243]. Society [493, 357, 23, 72, 120, 165, 198, 244, 286, 388, 491, 509, 632, 50, 87, 131, 177, 454, 25, 222]. Sociological [292, 449]. Sociologie [344]. Sociology [432]. Solvay [344]. Some [418, 485, 136, 588, 357, 135, 596, 123]. son [639]. Sources [247, 253]. sous [473]. South [95, 69]. Southern [565]. Southport [172]. Soziale [620]. Sozialstatistik [498]. Sozialstatistischen [620]. Sparkassen [327]. Special [573, 431, 224]. Spencer [370, 392]. Spirits [48]. Staaten [406, 585]. Staats [406]. Staats- [406]. Standard [236, 492, 663, 601, 547]. Standards [464]. Stande [638]. Standpoint [466]. Stanley [350]. State [467, 126, 572, 48, 60, 685]. Statement [341, 493, 291, 329]. Statements [247, 253]. States [309, 508, 684, 325, 400, 431, 520]. Statistica [602]. Statistical [11, 21, 35, 45, 23, 72, 57, 70, 84, 97, 108, 118, 128, 141, 120, 153, 163, 174, 184, 165, 193, 204, 219, 229, 198, 234, 246, 239, 249, 256, 268, 244, 279, 286, 280, 303, 333, 359, 419, 388, 382, 416, 451, 482, 491, 509, 504, 525, 556, 592, 653, 632, 626, 648, 671, 687, 111, 50, 87, 216, 254, 259, 655, 93, 131, 177, 467, 454, 232, 365, 25, 77, 227, 431, 317, 222, 513, 248, 636, 513]. Statistiches [429]. Statistician [466, 586]. Statistico [393]. Statistics [133, 92, 117, 162, 203, 213, 207, 208, 311, 678, 315, 264, 14, 16, 31, 53, 103, 236, 247, 253, 261, 306, 275, 527, 172, 515, 29, 49, 4, 113, 427, 77, 100, 196, 629, 156, 283, 486, 263, 252, 39, 17, 28, 41, 251, 396, 426, 469, 538, 551, 248, 363, 677, 515, 314, 463]. Statistik [343, 430, 620, 371, 574, 668, 637, 669]. Statistique [680, 316, 428, 639, 535, 621, 94, 680, 290, 535, 391, 514, 536]. Statistischen [343, 476, 603, 465]. Status [530, 587]. Statutory [251]. Sternegg [562]. Stock [139]. Strand [23, 72, 120, 165, 198, 244, 286, 388, 509, 632]. Streiks [574]. Strength [446]. Studien [585, 434]. Studies [317, 396, 586]. Studium [668]. Study [318, 330, 373, 564, 519, 495, 494, 431, 324, 354, 588, 622, 252, 123, 396, 464, 469, 551, 589]. Subsidies [78, 300]. Success [298]. Successive [91, 260]. such [232]. Sugar [122]. Suggested [655]. Suggestion [676]. Suggestions [232]. sui [569]. Suicide [431, 621]. Suisse [575]. sul [552]. Summarising [681]. Sundbarg [479]. Supérieur [391]. Supino [478]. Supplement [291, 4, 566, 568, 423]. Supply [468, 631, 405]. Surpeuplement [439]. Surplus [468]. Surproduction [374]. Surroundings [357]. Survey [447, 550]. Suspension [89, 104]. Sweated [402]. Swing [336]. Syndicats [615]. System [664, 179]. Systems [447].

T [617]. Table [214, 65, 507, 159]. Tables [657, 681, 425, 291, 566, 40, 91, 114, 136]. Tailoring [666]. Take [138]. Taking [297, 3]. Tariff [472]. Tariffs [580]. Taxation [508]. Taylor [330]. Teaching [587]. Technische [585]. Technischen [581]. Teil [668]. Temporary [346]. Temps [319]. Ten [341, 243, 76, 159]. Termin [582]. Termin-Handel [582]. Terrace [23, 72, 120, 165, 198, 244, 286, 388, 509, 632]. Terrains [662]. Textes [521]. Th [581, 618]. Their [14, 40, 357, 347, 516, 224, 274]. Theodore [410]. Théorie [604]. Theory [657, 342, 292, 313, 500]. Thereby [112]. Thereunder [197]. Third [311, 389, 388, 276, 208]. Thirty [233]. Thomas [665, 571]. Thompson [373]. Three [489]. Time [346]. Together [14, 533, 606, 572, 357, 492]. Tonnage [100]. Total [341]. Town [436]. Towne [317]. Towns [16, 492, 533, 606, 158]. Trade [213, 666, 247, 253, 291, 515, 553, 487, 30, 492, 533, 606, 188, 79, 644, 682, 20, 424, 167, 5, 635, 236, 515, 329, 580]. Trades [16, 31, 53, 682, 666, 613]. Trading [441, 377]. Traffic [196]. Traités [521]. Transition [373, 299]. Travail [521, 604, 614]. Travaux [344]. Treasury [542]. Treating [421]. Treatment [14]. Treizième [577]. Troisieme [473]. Trombert [575]. Tronnier [537]. Trust [397]. Trusts [615]. Tuberculosis [461, 420, 396, 551]. Turot [439]. Twelfth [463]. Two [276].

U.S [515]. Udny [342].

Ueberkapitalisation [517]. Uebersichten [430]. Ultimate [291]. Unconscious [262]. Unconsidered [485]. Underclothing [666]. Unearned [664]. Unemployed [573, 188, 5]. Unemployment [613, 641, 640, 611]. Unification [467]. Union [236, 247, 253, 188, 5]. United [493, 684, 291, 325, 400, 492, 682, 431, 520, 635, 418, 44, 96, 117, 209, 207, 208, 309, 16, 31, 53, 103, 236, 247, 253, 261, 275, 144, 338, 277, 631, 273, 233, 252, 508, 60, 540]. University [586]. Unpublished [463]. Unskilled [573]. Unterlagen [465].

Untersuchung [603, 476]. Upon [325, 89, 276]. Urban [689, 187]. Use [676, 28, 41]. Used [169]. Utilities [499].

V [313]. Vade [548]. Valuation [146]. Value [589]. Values [346]. Vargha [578]. Variables [276]. Variation [28, 41]. Variazioni [552]. Various [447, 135]. Velocity [654]. Veränderungen [638]. Veranstaltet [476]. Vereinigten [406]. Verification [655]. Verteilung [618]. Verteuerung [534]. VI [16]. Viallate [518]. Victor [575, 546]. Vie [544]. Vieillards [375, 609]. Vierter [668]. View [227, 571]. VII [31]. VIII [53]. Viscount [364]. Vital [252]. Vol [472]. Volksichte [537]. Volkswirthschaftliche [581]. Voltaire [615]. Volume [1, 37, 47, 86, 99, 130, 143, 176, 186, 221, 231, 258, 270, 335, 361, 453, 484, 558, 594, 673]. Volumes [291]. vom [476, 637]. Vorlesungen [343]. Vue [612].

W [369, 404, 685, 583, 454, 397, 579, 516, 328, 446, 405, 350, 611, 426]. W.C [198, 244, 286, 388, 509, 632]. Wage [427, 402, 464]. Wage-Earner [464]. Wage-Earners [427]. Wages [493, 16, 31, 53, 103, 236, 247, 253, 261, 275, 676, 64, 347, 113, 547, 157, 320, 492, 533, 606, 531, 212, 459, 572, 115, 69, 112, 601, 31, 53, 103]. Wahlreform [465]. Während [638]. Wales [423, 689, 157, 91, 159, 283, 206, 456, 308, 38, 90, 187, 457, 95, 69, 274, 568]. Walter [477]. Wampach [475]. War [542]. Water [630]. Waxweiler [344]. Wealth [111, 169, 338, 486, 352]. Weather [362]. Webb [523, 355, 523, 355]. wegen [476]. Weights [179]. Welfare [471]. Welsford [446]. Weltwirthschaft [407]. Wertzuwachssteuer [549]. Wesley [547]. West [103, 512, 494]. Western [674]. Westminster [322, 564]. Wheat [515, 42, 617]. Wheel [352]. which [138, 425, 489]. Whose [312]. Wileman [539]. William [318, 372, 293, 329, 448, 623]. Williams [665, 579]. Wilson [597, 170]. Wirtschaftskrisen [517]. Witney [292]. Wm [624, 431]. Wohnungen [476]. Wohnungsfrage [661]. Wohnungsinspektion [476]. Wohnungswesens [661]. Wolff [404]. Women [666, 463, 530, 215, 235, 629, 320]. Woodland [266]. Woods [345]. Wool [470, 124]. Woollen [103]. Words [248]. Work [666, 463, 467, 320, 572, 513]. Working [492, 533, 606, 663, 60]. Working-Class [533]. Workman [318]. Workmen [227, 197, 265]. Workpeople [682]. Works [11, 21, 35, 45, 57, 70, 84, 97, 108, 118, 128, 141, 153, 163, 174, 184, 193, 204, 219, 229, 239, 249, 256, 268, 342, 279]. World [75, 17, 372]. Worsted [103]. Wright [598].

X [236]. XI [247]. XII [680, 253]. XIII [261]. XIV [275]. XIX [395]. XX [409].

Year

[23, 72, 120, 165, 198, 244, 341, 286, 388, 509, 632, 666, 367, 197, 265, 201, 538, 539]. Year-Book [538, 539]. Years [341, 16, 31, 53, 103, 236, 247, 253, 261, 243, 157, 76, 159, 531, 233, 38]. Yermoloff [408]. Yield [341]. York [663, 471, 464, 306, 115]. Yorkshire [103]. Young [490]. Yule [342]. Yves [643, 443].

Zahlreicher

[407]. Zealand [572]. Zehnjährigen [534]. Zeitraum [534]. Zizek [603]. Zolla [614]. Zuckerindustrie [581]. zum [668, 537]. zur [434, 620, 584].

References

Anonymous:1900:VIa

[1] Anonymous. Volume information. Journal of the Royal Statistical Society, 63 (1):i–viii, March 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979720.

Rozenraad:1900:IMM

[2] Cornelius Rozenraad. The international money market. Journal of the Royal Statistical Society, 63(1):1–40, March 1900. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979721.

Baines:1900:CTL

[3] J. A. Baines. On census-taking and its limitations. Journal of the Royal Statistical Society, 63(1):41–71, March 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979722.

Edgeworth:1900:RSM

[4] F. Y. Edgeworth. On the representation of statistics by mathematical formulae (supplement). Journal of the Royal Statistical Society, 63(1):72–81, March 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979723.

Wood:1900:TUE

[5] George H. Wood. Trade union expenditure on unemployed benefits since 1860. Journal of the Royal Statistical Society, 63(1):81–92, March 1900. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979724.

Sauerbeck:1900:PC

[6] A. Sauerbeck. Prices of commodities in 1899. Journal of the Royal Statistical Society, 63(1):92–106, March 1900. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979725.

Anonymous:1900:CPQ

[7] Anonymous. The census of 1901 and the proposed quinquennial census. Journal of the Royal Statistical Society, 63(1): 107–111, March 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979726.

Anonymous:1900:CHR

[8] Anonymous. Commercial history and review of 1889. Journal of the Royal Statistical Society, 63(1):112–130, March 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979727.

Anonymous:1900:FLM

[9] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1899. Journal of the Royal Statistical Society, 63(1):131–135, March 1900. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979728.

Anonymous:1900:EL

[10] Anonymous. English literture in 1899. Journal of the Royal Statistical Society, 63(1):135–137, March 1900. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979729.

Anonymous:1900:NESa

[11] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 63(1):137– 146, March 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979730.

Anonymous:1900:QLAa

[12] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 63(1):147–159, March 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979731.

Anonymous:1900:PR

[13] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 63(1): 160–182, March 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979732.

Barrett:1900:TJO

[14] Rosa M. Barrett. The treatment of juvenile offenders: Together with statistics of their numbers. Journal of the Royal Statistical Society, 63(2):183–271, June 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979656.

Bence-Jones:1900:CAB

[15] H. Bence-Jones. On the consumption of alcoholic beverages. Journal of the Royal Statistical Society, 63(2):272–296, June 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979657.

Bowley:1900:SWUa

[16] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (Part VI). wages in the building trades. — English towns. Journal of the Royal Statistical Society, 63 (2):297–315, June 1900. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979658.

Schou:1900:SWF

[17] Vilhelm Schou. The statistics of the world’s foreign commerce. Journal of the Royal Statistical Society, 63(2):315–330, June 1900. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979659.

Anonymous:1900:RAR

[18] Anonymous. Report on the agricultural returns for 1899. Journal of the Royal Statistical Society, 63(2):330–335, June 1900. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979660.

Anonymous:1900:PC

[19] Anonymous. The proposed census in 1906. Journal of the Royal Statistical Society, 63(2):335–336, June 1900. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979661.

Moos:1900:GFT

[20] Ferdinand Moos. Germany’s foreign trade in 1899. Journal of the Royal Statistical Society, 63(2):337–340, June 1900. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979662.

Anonymous:1900:NESb

[21] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 63(2):340–350, June 1900. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979663.

Anonymous:1900:QLAb

[22] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 63(2):351–366, June 1900. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979664.

Anonymous:1900:RCF

[23] Anonymous. Report of the Council for the financial year ended 31st December, 1899, and for the sessional year ending 19th June, 1900, presented at the sixty-sixth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 19th of June, 1900. Journal of the Royal Statistical Society, 63(3):367–380, September 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979812.

Bateman:1900:PSS

[24] Alfred E. Bateman. Proceedings of the sixty-sixth Annual General Meeting. Journal of the Royal Statistical Society, 63(3):381–382, September 1900. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979813.

Fowler:1900:MFM

[25] Henry Hartley Fowler. Municipal finance and municipal enterprise. The annual address of the Right Hon. Sir Henry Hartley Fowler, G.C.S.I., M.P., President of the Royal Statistical Society, session 1899–1900. delivered 15th May, 1900. Journal of the Royal Statistical Society, 63(3):383–407, September 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979814.

Anonymous:1900:PM

[26] Anonymous. Proceedings on the 15th May, 1900. Journal of the Royal Statistical Society, 63(3):408–409, September 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979815.

Dilke:1900:DEE

[27] Charles W. Dilke. The defence expenditure of the empire. Journal of the Royal Statistical Society, 63(3):410– 432, September 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979816.

Sheppard:1900:UACa

[28] W. F. Sheppard. On the use of auxiliary curves in statistics of continuous variation. Journal of the Royal Statistical Society, 63(3):433–458, September 1900. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979817.

Craigie:1900:AES

[29] P. G. Craigie. Address to the Economic Science and Statistics Section of the British Association, held at Bradford, 1900. Journal of the Royal Statistical Society, 63(3):459–479, September 1900. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979818.

Flux:1900:PMF

[30] A. W. Flux. Price movements in the foreign trade of France. Journal of the Royal Statistical Society, 63(3):480– 484, September 1900. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979819.

Bowley:1900:SWUb

[31] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (Part VII). Wages in the building trades–contd. Scotland and Ireland. Journal of the Royal Statistical Society, 63(3):485–497, September 1900. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979820.

Anonymous:1900:PS

[32] Anonymous. Professor Sidgwick. Journal of the Royal Statistical Society, 63 (3):497–498, September 1900. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979821.

Anonymous:1900:SJB

[33] Anonymous. Sir John Bennet Lawes, Bart., F.R.S. Journal of the Royal Statistical Society, 63(3):498–499, September 1900. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979822.

Anonymous:1900:ARG

[34] Anonymous. Agricultural returns of Great Britain, 1900. Journal of the Royal Statistical Society, 63(3):499–500, September 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979823.

Anonymous:1900:NESc

[35] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 63(3):500–512, September 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979824.

Anonymous:1900:QLAc

[36] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 63(3):512–526, September 1900. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979825.

Anonymous:1900:VIb

[37] Anonymous. Volume information. Journal of the Royal Statistical Society, 63 (4):701–714, December 1900. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979909.

Welton:1900:DPE

[38] Thomas A. Welton. On the distribution of population in England and Wales, and its progress in the period of ninety years from 1801 to 1891. Journal of the Royal Statistical Society, 63(4):527– 595, December 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979910.

Rubin:1900:PBR

[39] Marcus Rubin. Population and birthrate, illustrated from historical statistics. Journal of the Royal Statistical Society, 63(4):596–625, December 1900. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979911.

Hayward:1900:LTT

[40] T. E. Hayward. On life-tables: Their construction and practical application. Journal of the Royal Statistical Society, 63(4):625–636, December 1900. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979912.

Sheppard:1900:UACb

[41] W. F. Sheppard. On the use of auxiliary curves in statistics of continuous variation. Journal of the Royal Statistical Society, 63(4):637–648, December 1900. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979913.

Hooker:1900:FPW

[42] R. H. Hooker. Farm prices of wheat and maize in America, 1870–99. Journal of the Royal Statistical Society, 63(4):648– 657, December 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979914.

Contributor:1900:PAC

[43] Colinal Contributor. Progress of the Australasian colonies. Journal of the Royal Statistical Society, 63(4):657– 667, December 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979915.

Anonymous:1900:ARU

[44] Anonymous. Agricultural returns of the United Kingdom for 1900. Journal of the Royal Statistical Society, 63(4):668– 670, December 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979916.

Anonymous:1900:NESd

[45] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 63(4):670– 683, December 1900. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979917.

Anonymous:1900:QLAd

[46] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 63(4):683–700, December 1900. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979918.

Anonymous:1901:VIa

[47] Anonymous. Volume information. Journal of the Royal Statistical Society, 64 (1):i–viii, March 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979919.

Raffalovich:1901:SMS

[48] Alexis Raffalovich. The state monopoly of spirits in Russia, and its influence on the prosperity of the population. Journal of the Royal Statistical Society, 64 (1):1–30, March 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979920.

Danvers:1901:RIS

[49] F. C. Danvers. A review of Indian statistics. Journal of the Royal Statistical Society, 64(1):31–72, March 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979921.

Avebury:1901:GMN

[50] Lord Avebury. The growth of municipal and national expenditure. The annual address of the Right Hon. Lord Avebury, F.R.S., D.C.L., LL.D., President of the Royal Statistical Society, session 1900–01. delivered 19th February, 1901. Journal of the Royal Statistical Society, 64(1):73–84, March 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979922.

Anonymous:1901:PF

[51] Anonymous. Proceedings on the 19th February, 1901. Journal of the Royal Statistical Society, 64(1):85–86, March 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979923.

Sauerbeck:1901:PC

[52] A. Sauerbeck. Prices of commodities in 1900. Journal of the Royal Statistical Society, 64(1):87–101, March 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979924.

Bowley:1901:SWU

[53] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (Part VIII). Wages in the building trades–concluded. London. Journal of the Royal Statistical Society, 64(1):102–112, March 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979925.

Anonymous:1901:CHR

[54] Anonymous. Commercial history and review of 1900. Journal of the Royal Statistical Society, 64(1):113–132, March 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979926.

Anonymous:1901:FLM

[55] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1900. Journal of the Royal Statistical Society, 64(1):133–138, March 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979927.

Anonymous:1901:EL

[56] Anonymous. English literature in 1900. Journal of the Royal Statistical Society, 64(1):138–140, March 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979928.

Anonymous:1901:NESa

[57] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 64(1):140– 152, March 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979929.

Anonymous:1901:QLAa

[58] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 64(1):152–165, March 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979930.

Anonymous:1901:PR

[59] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 64(1): 166–188, March 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979931.

Sykes:1901:RSM

[60] John F. J. Sykes. The results of state, municipal, and organized private action on the housing of the working classes in London and in other large cities in the United Kingdom. Journal of the Royal Statistical Society, 64(2):189–264, June 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979942.

Rabino:1901:ENP

[61] Joseph Rabino. An Economist’s notes on Persia. Journal of the Royal Statistical Society, 64(2):265–291, June 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979943.

Anonymous:1901:MMB

[62] Anonymous. M. Maurice Block. Journal of the Royal Statistical Society, 64(2): 292–293, June 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979944.

Waters:1901:MEM

[63] A. C. Waters. A method for estimating mean populations in the last intercensal period. Journal of the Royal Statistical Society, 64(2):293–298, June 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979945.

Brabrook:1901:WCI

[64] E. W. Brabrook. Wages in the cotton industry of Ghent, according to sex and age. Journal of the Royal Statistical Society, 64(2):298–299, June 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979946.

Bailey:1901:NTM

[65] A. H. Bailey. New table of mortality. Journal of the Royal Statistical Society, 64(2):300–301, June 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979947.

Anonymous:1901:RAR

[66] Anonymous. Report on the agricultural returns for 1900. Journal of the Royal Statistical Society, 64(2):301–306, June 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979948.

Anonymous:1901:FRC

[67] Anonymous. First results of the census, 1901. Journal of the Royal Statistical Society, 64(2):307–314, June 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979949.

Baines:1901:IC

[68] J. A. Baines. The Indian census. Journal of the Royal Statistical Society, 64(2): 314–327, June 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979950.

Wood:1901:CAW

[69] George H. Wood. Changes in average wages in New South Wales, 1823– 98. Journal of the Royal Statistical Society, 64(2):327–335, June 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979951.

Anonymous:1901:NESb

[70] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 64(2):336–345, June 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979952.

Anonymous:1901:QLAb

[71] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 64(2):346–362, June 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979953.

Anonymous:1901:FYE

[72] Anonymous. For the financial year ended 31st December, 1900, and for the sessional year ending 18th June, 1901, presented at the sixty-seventh Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 18th of June, 1901. Journal of the Royal Statistical Society, 64(3):363– 377, September 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979861.

Anonymous:1901:PSS

[73] Anonymous. Proceedings of the sixtyseventh Annual General Meeting. Journal of the Royal Statistical Society, 64 (3):378–379, September 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979862.

Branford:1901:CNR

[74] Victor Verasis Branford. On the calculation of national resources. Journal of the Royal Statistical Society, 64(3):380– 414, September 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979863.

Hooper:1901:RGP

[75] Wynnard Hooper. The recent gold production of the world. Journal of the Royal Statistical Society, 64(3):415– 433, September 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979864.

Hayward:1901:DMA

[76] T. E. Hayward. On the different methods available for calculating mean numbers of population for the ten calendar years most nearly corresponding to an intercensal period. Journal of the Royal Statistical Society, 64(3):434– 444, September 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979865.

Giffen:1901:IGS

[77] Robert Giffen. The importance of general statistical ideas. Address to the economic science and statistics section of the British Association, held at Glasgow, 1901. Journal of the Royal Statistical Society, 64(3):444–461, September 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979866.

Ginsburg:1901:SS

[78] Benedict William Ginsburg. Shipping subsidies. Journal of the Royal Statistical Society, 64(3):461–484, September 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979867.

Hooker:1901:CMR

[79] R. H. Hooker. Correlation of the marriage-rate with trade. Journal of the Royal Statistical Society, 64(3):485– 492, September 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979868.

Baines:1901:CN

[80] J. A. Baines. Census notes. Journal of the Royal Statistical Society, 64(3):493– 526, September 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979869.

Willcox:1901:NAC

[81] W. F. Willcox. Note on American census practice. Journal of the Royal Statistical Society, 64(3):526–529, September 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979870.

Anonymous:1901:ENP

[82] Anonymous. An Economist’s notes on Persia. Journal of the Royal Statistical Society, 64(3):530, September 1901. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979871.

Anonymous:1901:ARG

[83] Anonymous. Agricultural returns of Great Britain, 1901. Journal of the Royal Statistical Society, 64(3):531–532, September 1901. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979872.

Anonymous:1901:NESc

[84] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 64(3):533–542, September 1901. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979873.

Anonymous:1901:QLAc

[85] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 64(3):542–556, September 1901. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979874.

Anonymous:1901:VIb

[86] Anonymous. Volume information. Journal of the Royal Statistical Society, 64 (4):711–728, December 1901. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979837.

Avebury:1901:LIB

[87] Lord Avebury. Local and Imperial burdens. the inaugural address of the Right Hon. Lord Avebury, P.C., F.R.S., D.C.L., LL.D., President of the Royal Statistical Society session 1901–02. delivered 19th November, 1901. Journal of the Royal Statistical Society, 64(4):557– 572, December 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979838.

Anonymous:1901:PN

[88] Anonymous. Proceedings on the 19th November, 1901. Journal of the Royal Statistical Society, 64(4):573, December 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979839.

Hooker:1901:SBP

[89] R. H. Hooker. The suspension of the Berlin Produce Exchange and its effect upon corn prices. Journal of the Royal Statistical Society, 64(4):574– 613, December 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2979840. See correction [104].

Welton:1901:FDP

[90] Thomas A. Welton. On the facts disclosed in the preliminary report as to the population of England and Wales in 1901. Journal of the Royal Statistical Society, 64(4):614–636, December 1901. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979841.

Hayward:1901:SLT

[91] T. E. Hayward. A series of life-tables for England and Wales for each successive decennium from 1841–50 to 1881–90, calculated by an abbreviated method. Journal of the Royal Statistical Society, 64(4):636–641, December 1901. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979842.

Anonymous:1901:CSE

[92] Anonymous. A civil service examination in statistics. Journal of the Royal Statistical Society, 64(4):641–645, December 1901. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979843.

Craigie:1901:RPE

[93] P. G. Craigie. Report on the proceedings of the eighth Congress of the International Statistical Institute at Buda Pesth. Journal of the Royal Statistical Society, 64(4):645–649, December 1901. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979844.

deFoville:1901:SO

[94] Alfred de Foville. La statistique et l’opinion. Journal of the Royal Statistical Society, 64(4):649–661, December 1901. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2979845. See

comments [264].

Wood:1901:CAP

[95] George H. Wood. Changes in average prices in New South Wales, 1820–98. Journal of the Royal Statistical Society, 64(4):661–676, December 1901. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979846.

Anonymous:1901:ARU

[96] Anonymous. Agricultural returns of the United Kingdom for 1901. Journal of the Royal Statistical Society, 64(4):677– 679, December 1901. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979847.

Anonymous:1901:NESd

[97] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 64(4):680– 692, December 1901. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979848.

Anonymous:1901:QLAd

[98] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 64(4):692–709, December 1901. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979849.

Anonymous:1902:VIa

[99] Anonymous. Volume information. Journal of the Royal Statistical Society, 65 (1):i–viii, March 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979697.

Glover:1902:TSD

[100] John Glover. Tonnage statistics of the decade 1891–1900. Journal of the Royal Statistical Society, 65(1):1–46, March 1902. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979698.

Giffen:1902:FR

[101] Robert Giffen. A financial retrospect, 1861–1901. Journal of the Royal Statistical Society, 65(1):47–85, March 1902. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979699.

Sauerbeck:1902:PC

[102] A. Sauerbeck. Prices of commodities in 1901. Journal of the Royal Statistical Society, 65(1):86–101, March 1902. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979700.

Bowley:1902:SWU

[103] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (Part IX). Wages in the worsted and woollen manufactures of the West Riding of Yorkshire. Journal of the Royal Statistical Society, 65(1): 102–126, March 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979701.

Hooker:1902:SBP

[104] R. H. Hooker. The suspension of the Berlin Produce Exchange: A correction. Journal of the Royal Statistical Society, 65(1):126–127, March 1902. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979702. See [89].

Anonymous:1902:CHR

[105] Anonymous. Commercial history and review of 1901. Journal of the Royal Statistical Society, 65(1):127–148, March 1902. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979703.

Anonymous:1902:FLM

[106] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1901. Journal of the Royal Statistical Society, 65(1):148–152, March 1902. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979704.

Anonymous:1902:EL

[107] Anonymous. English literature in 1901. Journal of the Royal Statistical Society, 65(1):152–154, March 1902. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979705.

Anonymous:1902:NESa

[108] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 65(1):154– 169, March 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979706.

Anonymous:1902:QLAa

[109] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 65(1):169–185, March 1902. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979707.

Anonymous:1902:PR

[110] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 65(1): 186–208, March 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979708.

Atkinson:1902:SRI

[111] Fred. J. Atkinson. A statistical review of the income and wealth of British India. Journal of the Royal Statistical Society, 65(2):209–283, June 1902. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979665.

Wood:1902:FLC

[112] George Henry Wood. Factory legislation, considered with reference to the wages, &c., of the operatives protected thereby. Journal of the Royal Statistical Society, 65(2):284–324, June 1902. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979666.

Edgeworth:1902:MRS

[113] F. Y. Edgeworth. Methods of representing statistics of wages and other groups not fulfilling the normal law of error. Journal of the Royal Statistical Society, 65(2):325–354, June 1902. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979667.

Hayward:1902:NLT

[114] T. E. Hayward. Notes on life-tables. Journal of the Royal Statistical Society, 65(2):354–358, June 1902. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979668.

Rowntree:1902:WY

[115] Mr. B. S. Rowntree and A. L. Bowley. Wages in York in 1899. Journal of the Royal Statistical Society, 65(2):359–361, June 1902. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979669.

Anonymous:1902:RAR

[116] Anonymous. Report on the agricultural returns for 1901. Journal of the Royal Statistical Society, 65(2):362–366, June 1902. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979670.

Anonymous:1902:RCI

[117] Anonymous. Report of the committee to inquire into the statistics available as a basis for estimating the production and consumption of meat and milk in the United Kingdom. Journal of the Royal Statistical Society, 65(2):367–371, June 1902. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979671.

Anonymous:1902:NESb

[118] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 65(2):371–383, June 1902. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979672.

Anonymous:1902:QLAb

[119] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 65(2):384–400, June 1902. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979673.

Anonymous:1902:RCF

[120] Anonymous. Report of the Council for the financial year ended 31st December, 1901, and for the sessional year ending 17th June, 1902, presented at the sixty-eighth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 17th of June, 1902. Journal of the Royal Statistical Society, 65(3):401–416, September 1902. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339447.

Anonymous:1902:PSE

[121] Anonymous. Proceedings of the sixtyeighth Annual General Meeting. Journal of the Royal Statistical Society, 65(3): 417–418, September 1902. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339448.

Guyot:1902:SIC

[122] Yves Guyot. The sugar industry on the continent. Journal of the Royal Statistical Society, 65(3):419–446, September 1902. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339449.

Welton:1902:SSP

[123] T. A. Welton. A study of some portions of the census of London for 1901. Journal of the Royal Statistical Society, 65(3):447–502, September 1902. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339450.

Thompson:1902:WPG

[124] R. J. Thompson. Wool prices in Great Britain, 1883–1901. Journal of the Royal Statistical Society, 65(3):503– 513, September 1902. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339451.

Morison:1902:IPI

[125] T. Morison. The instability of prices in India before 1861. Journal of the Royal Statistical Society, 65(3):513– 525, September 1902. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339452.

Moos:1902:SGB

[126] F. Moos. State of the German banks. Journal of the Royal Statistical Society, 65(3):525–527, September 1902. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339453.

Anonymous:1902:ARG

[127] Anonymous. Agricultural returns of Great Britain, 1902. Journal of the Royal Statistical Society, 65(3):528–529, September 1902. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339454.

Anonymous:1902:NESc

[128] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 65(3):529–552, September 1902. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339455.

Anonymous:1902:QLAc

[129] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 65(3):552–570, September 1902. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339456.

Anonymous:1902:VIb

[130] Anonymous. Volume information. Journal of the Royal Statistical Society, 65 (4):743–755, December 1902. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339694.

Craigie:1902:IAM

[131] Patrick George Craigie. The inaugural address of Major Patrick George Craigie, C.B., President of the Royal Statistical Society, session 1902–03. delivered 18th November, 1902. Journal of the Royal Statistical Society, 65(4):571– 609, December 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339695.

Anonymous:1902:PN

[132] Anonymous. Proceedings on the 18th November, 1902. Journal of the Royal Statistical Society, 65(4):610– 612, December 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339696.

Acworth:1902:ERS

[133] W. M. Acworth. English railway statistics. Journal of the Royal Statistical Society, 65(4):613–664, December 1902. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339697. See notes

[150].

Anonymous:1902:SJD

[134] Anonymous. Sir Juland Danvers. Journal of the Royal Statistical Society, 65 (4):665–666, December 1902. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339698.

Rew:1902:MSE

[135] R. Henry Rew. Memorandum on some estimates made by various authorities of the production of meat and milk. Journal of the Royal Statistical Society, 65 (4):666–679, December 1902. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339699.

Hayward:1902:SAN

[136] T. E. Hayward. Some additional notes on a “short” method of constructing life-tables. Journal of the Royal Statistical Society, 65(4):680–684, December 1902. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339700.

Wood:1902:IRP

[137] George Henry Wood. The investigation of retail prices. Journal of the Royal Statistical Society, 65(4):685–690, December 1902. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339701.

Anonymous:1902:RCA

[138] Anonymous. The report of the committee appointed to inquire into the existing arrangements for the collection and custody of local records, and as to the further measures which it may be advisable to take for the purpose. Journal of the Royal Statistical Society, 65(4):691– 692, December 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339702.

Anonymous:1902:ARA

[139] Anonymous. Agricultural returns, 1902 (acreage and live stock). Journal of the Royal Statistical Society, 65(4):692– 697, December 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339703.

Chapman:1902:NAA

[140] S. Chapman. Note on the Association of American Railway Accounting Officers. Journal of the Royal Statistical Society, 65(4):698–702, December 1902. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339704.

Anonymous:1902:NESd

[141] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 65(4):702– 721, December 1902. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339705.

Anonymous:1902:QLAd

[142] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 65(4):721–741, December 1902. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339706.

Anonymous:1903:VIa

[143] Anonymous. Volume information. Journal of the Royal Statistical Society, 66 (1):i–viii, March 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339477.

Brassey:1903:FFG

[144] T. A. Brassey. The finance of Federal government for the United Kingdom. Journal of the Royal Statistical Society, 66(1):1–30, March 1903. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339478.

Baines:1903:CE

[145] J. A. Baines. A census of the empire.

Journal of the Royal Statistical Society, 66(1):31–70, March 1903. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339479.

Spensley:1903:LV

[146] J. Calvert Spensley. The London valuation. Journal of the Royal Statistical Society, 66(1):71–86, March 1903. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339480.

Sauerbeck:1903:PC

[147] A. Sauerbeck. Prices of commodities in 1902. Journal of the Royal Statistical Society, 66(1):87–102, March 1903. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339481.

Atkinson:1903:RPI

[148] Fred J. Atkinson. Rupee prices in India, 1898–1901. Journal of the Royal Statistical Society, 66(1):103–118, March 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339482.

Welton:1903:MC

[149] T. A. Welton. Memorandum on the census of 1901. Journal of the Royal Statistical Society, 66(1):118–122, March 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339483.

Price-Williams:1903:NMA

[150] R. Price-Williams. Notes on Mr. Acworth’s paper. Journal of the Royal Statistical Society, 66(1):122–126, March 1903. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2339484. See

[133].

Anonymous:1903:CHR

[151] Anonymous. Commercial history and review of 1902. Journal of the Royal Statistical Society, 66(1):127–147, March 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339485.

Anonymous:1903:EL

[152] Anonymous. English literature in 1902.

Journal of the Royal Statistical Society, 66(1):148–149, March 1903. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339486.

Anonymous:1903:NESa

[153] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 66(1):150– 171, March 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339487.

Anonymous:1903:QLAa

[154] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 66(1):171–189, March 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339488.

Anonymous:1903:PR

[155] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 66(1): 190–212, March 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339489.

King:1903:SI

[156] Bolton King. Statistics of Italy. Journal of the Royal Statistical Society, 66(2): 213–272, June 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339233.

Fox:1903:AWE

[157] A. Wilson Fox. Agricultural wages in England and Wales during the last fifty years. Journal of the Royal Statistical Society, 66(2):273–359, June 1903. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339234. See notes [170].

Welton:1903:MPO

[158] Thomas A. Welton. Memorandum on primary occupations in the principal English towns in 1901. Journal of the Royal Statistical Society, 66(2):360–365, June 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339235.

Hayward:1903:NLT

[159] T. E. Hayward. A new life-table for England and Wales, based on the census enumerations of 1891 and 1901 and on the recorded deaths for the ten years 1891–1900. Journal of the Royal Statistical Society, 66(2):366–372, June 1903. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339236.

Anonymous:1903:C

[160] Anonymous. The census of 1901. Journal of the Royal Statistical Society, 66 (2):373, June 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339237.

Anonymous:1903:FLM

[161] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1902. Journal of the Royal Statistical Society, 66(2):374–377, June 1903. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339238.

Anonymous:1903:ASR

[162] Anonymous. Agricultural statistics, 1902. report on the agricultural returns for Great Britain. Journal of the Royal Statistical Society, 66(2):377–381, June 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339239.

Anonymous:1903:NESb

[163] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 66(2):382–406, June 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339240.

Anonymous:1903:QLAb

[164] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 66(2):407–422, June 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339241.

Anonymous:1903:RCF

[165] Anonymous. Report of the Council for the financial year ended 31st December, 1902, and for the sessional year ending 16th June, 1903, presented at the sixty-ninth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 16th of June, 1903. Journal of the Royal Statistical Society, 66(3):423–436, September 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339409.

Anonymous:1903:PSN

[166] Anonymous. Proceedings of the sixtyninth Annual General Meeting. Journal of the Royal Statistical Society, 66(3): 437–438, September 1903. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339410.

Thomas:1903:GDO

[167] D. A. Thomas. The growth and direction of our foreign trade in coal during the last half century. Journal of the Royal Statistical Society, 66(3):439– 533, September 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339411.

Chance:1903:DLPa

[168] William Chance. A decade of London pauperism, 1891 to 1901. Journal of the Royal Statistical Society, 66(3):534– 581, September 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339412.

Giffen:1903:WEH

[169] Robert Giffen. The wealth of the empire, and how it should be used. Journal of the Royal Statistical Society, 66(3):582– 598, September 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339413.

Bowley:1903:NMW

[170] A. L. Bowley. Notes on Mr. Wilson Fox’s paper. Journal of the Royal Statistical Society, 66(3):598–601, September 1903. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2339414. See

[157].

Anonymous:1903:ARG

[171] Anonymous. Agricultural returns of Great Britain, 1903. Journal of the Royal Statistical Society, 66(3):601–602, September 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339415.

Brabrook:1903:AES

[172] Edward W. Brabrook. Address to the Economic Science and Statistics Section of the British Association for the Advancement of Science, held at Southport, 1903. Journal of the Royal Statistical Society, 66(3):603–616, September 1903. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339416.

Sauerbeck:1903:INP

[173] A. Sauerbeck. Index numbers of prices.

Journal of the Royal Statistical Society, 66(3):616–620, September 1903. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339417.

Anonymous:1903:NESc

[174] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 66(3):620–640, September 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339418.

Anonymous:1903:QLAc

[175] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 66(3):640–656, September 1903. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339419.

Anonymous:1903:VIb

[176] Anonymous. Volume information. Journal of the Royal Statistical Society, 66 (4):779–792, December 1903. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339490.

Craigie:1903:AAM

[177] Patrick George Craigie. The annual address of Major Patrick George Craigie, C.B., President of the Royal Statistical Society, session 1903–04. delivered 17th November, 1903. Journal of the Royal Statistical Society, 66(4):657– 682, December 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339491.

Anonymous:1903:PN

[178] Anonymous. Proceedings on the 17th November, 1903. Journal of the Royal Statistical Society, 66(4):683– 687, December 1903. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339492.

Siemens:1903:MSW

[179] Alex. Siemens. The metrical system of weights and measures. Journal of the Royal Statistical Society, 66(4):688– 719, December 1903. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339493.

Ginsburg:1903:GR

[180] Benedict William Ginsburg. The growth of rates. Journal of the Royal Statistical Society, 66(4):720–728, December 1903. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339494.

Mackenzie:1903:IFC

[181] John A. P. Mackenzie. Investments of French capital abroad. Journal of the Royal Statistical Society, 66(4):729– 731, December 1903. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339495.

Chance:1903:DLPb

[182] William Chance. A decade of London pauperism. Journal of the Royal Statistical Society, 66(4):731–734, December 1903. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339496.

Anonymous:1903:AR

[183] Anonymous. Agricultural returns, 1903. Journal of the Royal Statistical Society, 66(4):735–740, December 1903. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339497.

Anonymous:1903:NESd

[184] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 66(4):740– 762, December 1903. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339498.

Anonymous:1903:QLAd

[185] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 66(4):762–778, December 1903. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339499.

Anonymous:1904:VIa

[186] Anonymous. Volume information. Journal of the Royal Statistical Society, 67 (1):i–viii, March 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339333.

Welton:1904:SUD

[187] Thomas A. Welton. On the smaller urban districts of England and Wales. Journal of the Royal Statistical Society, 67(1):1–51, March 1904. ISSN 0952-

8385. URL http://www.jstor.org/

stable/2339334. See errata [689].

Hartley:1904:TUE

[188] E. L. Hartley. Trade union expenditure on unemployed benefit. Journal of the Royal Statistical Society, 67(1): 52–80, March 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339335.

Schooling:1904:OEC

[189] John Holt Schooling. Our exports of coal-capital. Journal of the Royal Statistical Society, 67(1):81–85, March 1904. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339336.

Sauerbeck:1904:PC

[190] A. Sauerbeck. Prices of commodities in 1903. Journal of the Royal Statistical Society, 67(1):85–99, March 1904. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339337.

Anonymous:1904:CHR

[191] Anonymous. Commercial history and review of 1903. Journal of the Royal Statistical Society, 67(1):100–121, March 1904. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339338.

Anonymous:1904:EL

[192] Anonymous. English literature in 1903. Journal of the Royal Statistical Society, 67(1):122–123, March 1904. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339339.

Anonymous:1904:NESa

[193] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 67(1):123– 140, March 1904. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339340.

Anonymous:1904:QLAa

[194] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 67(1):140–155, March 1904. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339341.

Anonymous:1904:PR

[195] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 67(1):

156–176, March 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339342.

Harper:1904:SLT

[196] Edgar J. Harper. Statistics of London traffic. Journal of the Royal Statistical Society, 67(2):177–227, June 1904. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339597.

Tozer:1904:FYE

[197] William H. Tozer. Five years’ experience of the effect of the Workmen’s Compensation Acts, with especial reference to schemes certified thereunder. Journal of the Royal Statistical Society, 67(2): 228–256, June 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339598.

Anonymous:1904:RCF

[198] Anonymous. Report of the Council for the financial year ended 31st December, 1903, and for the sessional year ending 21st June, 1904, presented at the seventieth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, W.C., London, on the 21st of June, 1904. Journal of the Royal Statistical Society, 67(2):257–267, June 1904. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339599.

Anonymous:1904:PSA

[199] Anonymous. Proceedings of the seventieth Annual General Meeting. Journal of the Royal Statistical Society, 67(2): 268–271, June 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339600.

Anonymous:1904:PDP

[200] Anonymous. Proceedings of the deputation to the President of the Local Government Board, Friday, 13th May, 1904. Journal of the Royal Statistical Society, 67(2):272–281, June 1904. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339601.

Welton:1904:NRG

[201] Thomas A. Welton. Note on the Registrar-General’s Sixty-Fourth Annual Report, being that for the year 1901. Journal of the Royal Statistical Society, 67(2):281–284, June 1904. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339602.

Anonymous:1904:FLM

[202] Anonymous. Fires in London and the Metropolitan Fire Brigade in 1903. Journal of the Royal Statistical Society, 67(2):284–287, June 1904. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339603.

Anonymous:1904:ASR

[203] Anonymous. Agricultural statistics, 1903. report on the agricultural returns for Great Britain. Journal of the Royal Statistical Society, 67(2):287–290, June 1904. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339604.

Anonymous:1904:NESb

[204] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 67(2):290–312, June 1904. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339605.

Anonymous:1904:QLAb

[205] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 67(2):312–336, June 1904. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339606.

Thompson:1904:LEL

[206] R. J. Thompson. Local expenditure and local indebtedness in England and Wales. Journal of the Royal Statistical Society, 67(3):337–367, September 1904. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339432.

Anonymous:1904:PCMa

[207] Anonymous. Production and consumption of meat and milk. Second report from the committee appointed to inquire into the statistics available as a basis for estimating the production and consumption of meat and milk in the United Kingdom. Journal of the Royal Statistical Society, 67(3):368–384, September 1904. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339433.

Anonymous:1904:PCMb

[208] Anonymous. Production and consumption of meat and milk. Third report from the committee appointed to inquire into the statistics available as a basis for estimating the production and consumption of meat and milk in the United Kingdom. Journal of the Royal Statistical Society, 67(3):385–393, September 1904. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339434.

Anonymous:1904:ARP

[209] Anonymous. Appendix to the reports on the production and consumption of meat and milk in the United Kingdom. Journal of the Royal Statistical Society, 67(3):394–412, September 1904. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339435.

Rew:1904:OPC

[210] R. H. Rew. Observations on the production and consumption of meat and dairy products. Journal of the Royal Statistical Society, 67(3):413–427, September 1904. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339436.

Anonymous:1904:PMJ

[211] Anonymous. Proceedings of the meeting on 21st June, 1904. Journal of the Royal Statistical Society, 67(3):427– 429, September 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339437.

Mahaim:1904:CWR

[212] E. Mahaim. Changes in wages, and real wages in Belgium. Journal of the Royal Statistical Society, 67(3):430– 438, September 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339438.

Anonymous:1904:ERC

[213] Anonymous. Extracts from the report of the committee of the British Association appointed to inquire into the accuracy and comparability of British and foreign statistics of international trade. Journal of the Royal Statistical Society, 67(3):438–447, September 1904. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339439.

Adam:1904:LTS

[214] T. Adam. Life-table for Scotland, based on the census enumerations of 1891 and 1901, and on the recorded deaths for the decennium 1891–1900. Journal of the Royal Statistical Society, 67(3):448– 478, September 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339440.

Hutchins:1904:NDW

[215] B. L. Hutchins. A note on the distribution of women in occupations. Journal of the Royal Statistical Society, 67(3):479– 490, September 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339441.

Baines:1904:ISI

[216] J. A. Baines. The International Statistical Institute. Journal of the Royal Statistical Society, 67(3):490–497, September 1904. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339442.

Dudfield:1904:FNA

[217] Reginald Dudfield. A further note on the annual report of the Registrar-General. Journal of the Royal Statistical Society, 67(3):498–499, September 1904. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339443.

Anonymous:1904:ARGa

[218] Anonymous. Agricultural returns of Great Britain, 1904. Journal of the Royal Statistical Society, 67(3):500–501, September 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339444.

Anonymous:1904:NESc

[219] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 67(3):501–541, September 1904. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339445.

Anonymous:1904:QLAc

[220] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 67(3):541–560, September 1904. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339446.

Anonymous:1904:VIb

[221] Anonymous. Volume information. Journal of the Royal Statistical Society, 67 (4):749–761, December 1904. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339651.

Powell:1904:PAS

[222] Francis Sharp Powell. Presidential address of Sir Francis Sharp Powell, Bart., M.P., delivered to the Royal Statistical Society, 15th November, 1904. Journal of the Royal Statistical Society, 67(4): 561–581, December 1904. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339652.

Anonymous:1904:PN

[223] Anonymous. Proceedings on the 15th November, 1904. Journal of the Royal Statistical Society, 67(4):582– 584, December 1904. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339653.

Price:1904:ACO

[224] L. L. Price. The accounts of the Colleges of Oxford, 1893–1903; with special reference to their agricultural revenues. Journal of the Royal Statistical Society, 67(4):585–660, December 1904. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339654.

Baines:1904:GDP

[225] J. A. Baines. The geographical distribution of population in India. Journal of the Royal Statistical Society, 67(4):661– 670, December 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339655.

Fehlinger:1904:ACO

[226] H. Fehlinger. The Austrian census of occupations, 1900. Journal of the Royal Statistical Society, 67(4):671– 675, December 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339656.

Gladwell:1904:SVW

[227] S. W. Gladwell. A statistical view of the Workmen’s Compensation Act Committee’s report. Journal of the Royal Statistical Society, 67(4):675–678, December 1904. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339657.

Anonymous:1904:ARGb

[228] Anonymous. Agricultural returns of Great Britain, 1904. Journal of the Royal Statistical Society, 67(4):679– 683, December 1904. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339658.

Anonymous:1904:NESd

[229] Anonymous. Notes on economic and statistical works. Journal of the Royal Statistical Society, 67(4):684–732, December 1904. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339659.

Anonymous:1904:QLAd

[230] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 67(4):732–748, December 1904. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339660.

Anonymous:1905:VIa

[231] Anonymous. Volume information. Journal of the Royal Statistical Society, 68 (1):i–viii, March 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339322.

Dudfield:1905:CEM

[232] Reginald Dudfield. A critical examination of the methods of recording and publishing statistical data bearing on public health; with suggestions for the improvement of such methods. Journal of the Royal Statistical Society, 68 (1):1–49, March 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339323.

Palgrave:1905:EAL

[233] R. H. Inglis Palgrave. Estimates of agricultural losses in the United Kingdom during the last thirty years. Journal of the Royal Statistical Society, 68(1): 50–92, March 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339324.

Anonymous:1905:ISIa

[234] Anonymous. International Statistical Institute. London session, 31st July to 5th August, 1905. Journal of the Royal Statistical Society, 68(1):93–94, March 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339325.

Hutchins:1905:NDM

[235] B. L. Hutchins. Note on the distribution of married women in relation to the birth-rate. Journal of the Royal Statistical Society, 68(1):95–103, March 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339326.

Bowley:1905:SWUa

[236] A. L. Bowley and George H. Wood.

The statistics of wages in the United Kingdom during the last hundred years (Part X.) Engineering and shipbuilding. A. Trade union standard rates. Journal of the Royal Statistical Society, 68(1): 104–137, March 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339327.

Sauerbeck:1905:PC

[237] A. Sauerbeck. Prices of commodities in 1904. Journal of the Royal Statistical Society, 68(1):138–153, March 1905. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339328.

Anonymous:1905:CHR

[238] Anonymous. Commercial history and review of 1904. Journal of the Royal Statistical Society, 68(1):154–175, March 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339329.

Anonymous:1905:NESa

[239] Anonymous. Notes on economic and statistical works. Journal of the Royal Statistical Society, 68(1):176–208, March 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339330.

Anonymous:1905:QLAa

[240] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 68(1):209–227, March 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339331.

Anonymous:1905:PR

[241] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 68(1): 228–246, March 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339332.

Shaw:1905:SBI

[242] W. N. Shaw. Seasons in the British Isles from 1878. Journal of the Royal Statistical Society, 68(2):247–319, June 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339607.

Brabrook:1905:PFS

[243] E. W. Brabrook. On the progress of Friendly Societies and other institutions connected with the Friendly Societies Registry Office during the ten years 1894–1904. Journal of the Royal Statistical Society, 68(2):320–352, June 1905. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339608.

Anonymous:1905:RCF

[244] Anonymous. Report of the Council for the financial year ended 31st December, 1904, and for the sessional year ending 27th June, 1905, presented at the seventy-first Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, W.C., London, on the 27th of June, 1905. Journal of the Royal Statistical Society, 68(2):353–363, June 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339609.

Anonymous:1905:PSF

[245] Anonymous. Proceedings of the seventyfirst Annual General Meeting. Journal of the Royal Statistical Society, 68(2):

364–365, June 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339610.

Anonymous:1905:ISIb

[246] Anonymous. International Statistical Institute, London session, 31st July to 4th August, 1905. Journal of the Royal Statistical Society, 68(2):366–373, June 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339611.

Bowley:1905:SWUb

[247] A. L. Bowley. The statistics of wages in the United Kingdom during the last hundred years. (Part XI). Engineering and shipbuilding. B. Statements of wages from non-trade union sources in general engineering. Journal of the Royal Statistical Society, 68(2):373–391, June 1905. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339612.

Yule:1905:IWS

[248] G. Udny Yule. The introduction of the words “statistics,” “statistical” into the English language. Journal of the Royal Statistical Society, 68(2):391–396, June 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339613.

Anonymous:1905:NESb

[249] Anonymous. Notes on economic and statistical works. Journal of the Royal Statistical Society, 68(2):396–416, June 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339614.

Anonymous:1905:QLAb

[250] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 68(2):416–434, June 1905. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339615.

Ward:1905:ESS

[251] Leonard Ward. The effect, as shown by statistics, of British statutory regulations directed to the improvement of the hygienic conditions of industrial occupations. Journal of the Royal Statistical Society, 68(3):435–525, September 1905. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339590.

Rosenbaum:1905:CSV

[252] S. Rosenbaum. A contribution to the study of the vital and other statistics of the Jews in the United Kingdom. Journal of the Royal Statistical Society, 68(3):526–562, September 1905. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339591.

Bowley:1905:SWUc

[253] A. L. Bowley and George H. Wood. The statistics of wages in the United Kingdom during the last hundred years. (Part XII). Engineering and shipbuilding. C. Statements of wages from non-trade union sources in shipbuilding and engineering at shipbuilding centres. Journal of the Royal Statistical Society, 68(3): 563–614, September 1905. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339592.

Baines:1905:ISI

[254] J. Athelstane Baines. The International Statistical Institute. Journal of the Royal Statistical Society, 68(3):614– 625, September 1905. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339593.

Anonymous:1905:ARGa

[255] Anonymous. Agricultural returns of Great Britain, 1905. Journal of the Royal Statistical Society, 68(3):626–627, September 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339594.

Anonymous:1905:NESc

[256] Anonymous. Notes on economic and statistical works. Journal of the Royal Statistical Society, 68(3):627–655, September 1905. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339595.

Anonymous:1905:QLAc

[257] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 68(3):656–672, September 1905. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339596.

Anonymous:1905:VIb

[258] Anonymous. Volume information. Journal of the Royal Statistical Society, 68 (4):789–799, December 1905. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339420.

Baines:1905:SSI

[259] J. Athelstane Baines. Statistical skimmings from the International Congress. Journal of the Royal Statistical Society, 68(4):673–695, December 1905. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339421.

Hooker:1905:CSO

[260] R. H. Hooker. On the correlation of successive observations. Journal of the Royal Statistical Society, 68(4):696– 703, December 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339422.

Bowley:1905:SWUd

[261] A. L. Bowley and George H. Wood. The statistics of wages in the United Kingdom during the last hundred years. (Part XIII). Engineering and shipbuilding. D. Dockyards and railway centres. Journal of the Royal Statistical Society, 68(4): 704–715, December 1905. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339423.

Cunningham:1905:UAE

[262] W. Cunningham. Unconscious assumptions in economics. Journal of the Royal Statistical Society, 68(4):715– 724, December 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339424.

Mandello:1905:FS

[263] J. G. Mandello. The future of statistics. Journal of the Royal Statistical Society, 68(4):725–732, December 1905. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339425.

Baines:1905:MFS

[264] J. Athelstane Baines. M. de Foville on statistics. Journal of the Royal Statistical Society, 68(4):733–734, December 1905. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2339426. See

[94].

Tozer:1905:CCI

[265] William H. Tozer. Completed cases of incapacity from injury during the year ending 30th June, 1904, deduced from the annual returns of certified schemes under the Workmen’s Compensation Acts, 1897 and 1900. Journal of the Royal Statistical Society, 68(4):735– 737, December 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339427.

Anonymous:1905:WAG

[266] Anonymous. Woodland area of Great Britain, 1905. Journal of the Royal Statistical Society, 68(4):737–739, December 1905. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339428.

Anonymous:1905:ARGb

[267] Anonymous. Agricultural returns of Great Britain, 1905. Journal of the Royal Statistical Society, 68(4):739– 744, December 1905. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339429.

Anonymous:1905:NESd

[268] Anonymous. Notes on economic and statistical works. Journal of the Royal Statistical Society, 68(4):745–768, December 1905. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339430.

Anonymous:1905:QLAd

[269] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 68(4):768–787, December 1905. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339431.

Anonymous:1906:VIa

[270] Anonymous. Volume information. Journal of the Royal Statistical Society, 69 (1):i–viii, March 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339546.

ofOnslow:1906:RDL

[271] Earl of Onslow. The rise and development of local legislation by private bill. The Presidential address of the Right Hon. The Earl of Onslow, G.C.M.G. delivered 16th January, 1906. Journal of the Royal Statistical Society, 69 (1):1–31, March 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339547.

Anonymous:1906:PJ

[272] Anonymous. Proceedings on January 16th, 1906. Journal of the Royal Statistical Society, 69(1):32–33, March 1906. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339548.

Newsholme:1906:DHF

[273] Arthur Newsholme and T. H. C. Stevenson. The decline of human fertility in the United Kingdom and other countries as shown by corrected birth-rates. Journal of the Royal Statistical Society, 69 (1):34–87, March 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339549.

Yule:1906:CMB

[274] G. Udny Yule. On the changes in the marriage- and birth-rates in England and Wales during the past half century; with an inquiry as to their probable causes. Journal of the Royal Statistical Society, 69(1):88–147, March 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339550.

Bowley:1906:SWU

[275] A. L. Bowley and George H. Wood. The statistics of wages in the United Kingdom during the Nineteenth Century (Part XIV). Engineering and shipbuilding. E. Averages, index numbers, and general results. Journal of the Royal Statistical Society, 69(1):148–196, March 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339551.

Hooker:1906:NER

[276] R. H. Hooker and G. U. Yule. Note on estimating the relative influence of two variables upon a third. Journal of the Royal Statistical Society, 69(1): 197–200, March 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339552.

Haynes:1906:AAI

[277] Thomas H. Haynes. An analysis of Australian imports of merchandise from the United Kingdom and foreign countries. Journal of the Royal Statistical Society, 69(1):200–211, March 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339553.

Sauerbeck:1906:PC

[278] A. Sauerbeck. Prices of commodities in 1905. Journal of the Royal Statistical Society, 69(1):211–226, March 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339554.

Anonymous:1906:NES

[279] Anonymous. Notes on economic and statistical works. Journal of the Royal Statistical Society, 69(1):227–246, March 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339555.

Anonymous:1906:SEAa

[280] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 69(1): 247–253, March 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339556.

Anonymous:1906:QLAa

[281] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 69(1):254–269, March 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339557.

Anonymous:1906:PR

[282] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 69(1): 270–288, March 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339558.

Loch:1906:SPP

[283] C. S. Loch. Statistics of population and pauperism in England and Wales, 1861– 1901. Journal of the Royal Statistical Society, 69(2):289–320, June 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339524.

Chapman:1906:DFC

[284] S. J. Chapman and Douglas Knoop. Dealings in futures on the cotton market. Journal of the Royal Statistical Society, 69(2):321–373, June 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339525.

Thompson:1906:DAD

[285] R. J. Thompson. The development of agriculture in Denmark. Journal of the Royal Statistical Society, 69(2):374–419, June 1906. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339526.

Anonymous:1906:RCF

[286] Anonymous. Report of the Council for the financial year ended 31st December, 1905, and for the sessional year ending 19th June, 1906, presented at the seventy-second Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, W.C., London, on the 19th of June, 1906. Journal of the Royal Statistical Society, 69(2):420–430, June 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339527.

Anonymous:1906:PSS

[287] Anonymous. Proceedings of the seventysecond Annual General Meeting. Journal of the Royal Statistical Society, 69 (2):431–432, June 1906. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339528.

Flux:1906:CCR

[288] A. W. Flux. Canadian census results.

Journal of the Royal Statistical Society, 69(2):433–440, June 1906. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339529.

Baines:1906:PBE

[289] J. Athelstane Baines. The population of the British empire. Journal of the Royal Statistical Society, 69(2):440–443, June 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339530.

Y:1906:BRBa

[290] G. U. Y. Book review: Elements de Statistique, by Fernand Faure. Journal of the Royal Statistical Society, 69(2): 444–445, June 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339531.

C:1906:BRBa

[291] W. H. C. Book review: Abstract and Detailed Tables Showing Countries of Consignment of Imports and Countries of Ultimate Destination of Exports, with a Prefatory Memorandum, Being a Supplement to Volumes I and II of the Annual Statement of Trade of the United Kingdom for 1904. Journal of the Royal Statistical Society, 69(2):446–448, June 1906. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339532.

P:1906:BRBa

[292] L. L. P. Book review: The Sociological Theory of Capital; Being a Complete Reprint of the New Principles of Political Economy, 1834, by John Rae; Charles Witney Mixter. Journal of the Royal Statistical Society, 69(2):448–450, June 1906. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339533.

P:1906:BRBb

[293] L. L. P. Book review: Foundations of Political Economy, by William Bell Robertson. Journal of the Royal Statistical Society, 69(2):450–453, June 1906. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339534.

N:1906:BRBa

[294] A. N. Book review: Rapport sur les Causes Economiques de la Depopulation, by M. Alfred Neymarck. Journal of the Royal Statistical Society, 69(2): 453–454, June 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339535.

N:1906:BRBb

[295] A. N. Book review: Rapport sur les Causes Professionnelles de Depopulation, by M. Lucien March. Journal of the Royal Statistical Society, 69(2):454–455, June 1906. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339536.

B:1906:BRBa

[296] A. L. B. Book review: Des Retenues sur les Salaires Operees Pour le Compte du Patron, by Maurice Igier. Journal of the Royal Statistical Society, 69(2):455–457, June 1906. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339537.

B:1906:BRBb

[297] E. B. Book review: A History of English Philanthropy from the Dissolution of the Monasteries to the Taking of the First Census, by B. Kirkman Gray. Journal of the Royal Statistical Society, 69(2): 457–460, June 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339538.

P:1906:BRBc

[298] L. L. P. Book review: Modern Germany: Her Political and Economic Problems, Her Policy, Her Ambitions, and the Causes of Her Success, by O. Eltzbacher. Journal of the Royal Statistical Society, 69(2):460–463, June 1906. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339539.

G:1906:BRBa

[299] E. H. G. Book review: The Transition in Agriculture, by Edwin A. Pratt. Journal of the Royal Statistical Society, 69(2): 463–466, June 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339540.

G:1906:BRBb

[300] B. W. G. Book review: History of Shipping Subsidies, by Royal Meeker. Journal of the Royal Statistical Society, 69 (2):466–468, June 1906. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339541.

Anonymous:1906:C

[301] Anonymous. Correspondence. Journal of the Royal Statistical Society, 69(2):

468–469, June 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339542.

Anonymous:1906:ONPa

[302] Anonymous. Other new publications. Journal of the Royal Statistical Society, 69(2):470–471, June 1906. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339543.

Anonymous:1906:SEAb

[303] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 69(2): 472–480, June 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339544.

Anonymous:1906:QLAb

[304] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 69(2):481–496, June 1906. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339545.

Edgeworth:1906:GLE

[305] F. Y. Edgeworth. The Generalised Law of Error, or Law of Great Numbers. Journal of the Royal Statistical Society, 69(3):497–539, September 1906. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339343.

Bowley:1906:AES

[306] A. L. Bowley. Address to the economic science and statistics section of the British Association for the Advancement of Science, York, 1906. Journal of the Royal Statistical Society, 69(3):540– 558, September 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339344.

Foldes:1906:C

[307] Bela Foldes. The criminal. Journal of the Royal Statistical Society, 69(3):558– 576, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339345.

Vigor:1906:SRB

[308] H. D. Vigor and G. Udny Yule. On the sex-ratios of births in the registration districts of England and Wales, 1881–90. Journal of the Royal Statistical Society, 69(3):576–582, September 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339346.

Anonymous:1906:USC

[309] Anonymous. United States. Change in classification of imports and exports. Journal of the Royal Statistical Society, 69(3):582–585, September 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339347.

Anonymous:1906:ARG

[310] Anonymous. Agricultural returns of Great Britain, 1906. Journal of the Royal Statistical Society, 69(3):586–587, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339348.

Anonymous:1906:TAF

[311] Anonymous. Third abstract of foreign labour statistics. Journal of the Royal Statistical Society, 69(3):588–590, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339349.

Y:1906:BRBb

[312] G. U. Y. Book review: Noteworthy Families (Modern Science): An Index to Kinships in Near Degrees Between Persons Whose Achievements are Honourable, and Have been Publicly Recorded, by Francis Galton; Edgar Schuster. Journal of the Royal Statistical Society, 69(3):590–593, September 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339350.

Y:1906:BRBc

[313] G. U. Y. Book review: Meddelanden Fran Lunds Astronomiska Observatorium. Serie II, Nr. 4. Researches into the Theory of Probability, by C. V. L. Charlier. Journal of the Royal Statistical Society, 69(3):593, September 1906. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339351.

Y:1906:BRBd

[314] G. U. Y. Book review: A Brief Introduction to the Infinitesimal Calculus; Designed Especially to Aid in Reading Mathematical Economics and Statistics, by Irving Fisher. Journal of the Royal Statistical Society, 69(3):593– 594, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339352.

B:1906:BRBc

[315] J. A. B. Book review: London Statistics, 1905–06. Journal of the Royal Statistical Society, 69(3):594–596, September 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339353.

B:1906:BRBd

[316] A. E. B. Book review: Apercu des Resultats de la Statistique Comparee du Commerce Exterieur, by E. Levasseur. Journal of the Royal Statistical Society, 69(3):596–598, September 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339354.

P:1906:BRBd

[317] L. L. P. Book review: Economic and Statistical Studies, 1840–90, by John Towne; Mary Norman Hill; E. C. K. Gonner. Journal of the Royal Statistical Society, 69(3):598–601, September 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339355.

B:1906:BRBe

[318] A. L. B. Book review: The German Workman: A Study in National Efficiency, by William Harbutt Dawson. Journal of the Royal Statistical Society, 69(3):601–603, September 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339356.

B:1906:BRBf

[319] A. L. B. Book review: Ouvriers du Temps Passe (15^{e} –16^{e} Siecles), by H. Hauser. Journal of the Royal Statistical Society, 69(3):603–604, September 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339357.

H:1906:BRBa

[320] B. L. H. Book review: Women’s Work and Wages: A Phase of Life in an Industrial City, by Edw. Cadbury; M. C. Matheson; G. Shann. Journal of the Royal Statistical Society, 69(3):604– 606, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339358.

P:1906:BRBe

[321] L. L. P. Book review: La Revolution Industrielle au 18^{e} Siecle: Essai sur les Commencements de la Grande Industrie Moderne en Angleterre, by Paul Mantoux. Journal of the Royal Statistical Society, 69(3):606–608, September 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339359.

N:1906:BRBc

[322] A. N. Book review: National Conference on Infantile Mortality. Report of the Proceedings of the National Conference on Infantile Mortality Held in the Caxton Hall, Westminster, on the 13th and 14th June, 1906. Journal of the Royal Statistical Society, 69(3):608– 609, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339360.

N:1906:BRBd

[323] A. N. Book review: Infant Mortality; A Social Problem, by George Newman. Journal of the Royal Statistical Society, 69(3):610–611, September 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339361.

P:1906:BRBf

[324] L. L. P. Book review: Industrial Efficiency: A Comparative Study of Industrial Life in England, Germany, and America, by Arthur Shadwell. Journal of the Royal Statistical Society, 69(3): 611–614, September 1906. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339362.

C:1906:BRBb

[325] W. H. C. Book review: Immigration and its Effects Upon the United States, by Prescott F. Hall. Journal of the Royal Statistical Society, 69(3):614– 616, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339363.

P:1906:BRBg

[326] R. H. I. P. Book review: The Principles of Money and Banking, by Charles A. Conant. Journal of the Royal Statistical Society, 69(3):617–618, September 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339364.

B:1906:BRBg

[327] E. B. Book review: Die Deutschen Sparkassen in Bohmen, by Heinrich Rauchberg; K. K. Deutschen. Journal of the Royal Statistical Society, 69(3):618– 620, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339365.

P:1906:BRBh

[328] L. L. P. Book review: Retaliatory Duties, by H. Dietzel; D. W. Simon; W. Osborne Brigstocke. Journal of the Royal Statistical Society, 69(3):621– 622, September 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339366.

P:1906:BRBi

[329] L. L. P. Book review: The Return to Protection: Being a Re-Statement of the Case for Free Trade, by William Smart. Journal of the Royal Statistical Society, 69(3):623–625, September 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339367.

C:1906:BRBc

[330] P. G. C. Book review: An Introduction to the Study of Agricultural Economics, by Henry C. Taylor. Journal of the Royal Statistical Society, 69(3):625– 627, September 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339368.

F:1906:BRB

[331] A. W. F. Book review: The Canadian Annual Review of Public Affairs, 1905, by J. Castell Hopkins. Journal of the Royal Statistical Society, 69(3):627– 628, September 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339369.

Anonymous:1906:ONPb

[332] Anonymous. Other new publications. Journal of the Royal Statistical Society, 69(3):628–630, September 1906. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339370.

Anonymous:1906:SEAc

[333] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 69(3):630– 637, September 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339371.

Anonymous:1906:QLAc

[334] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 69(3):638–654, September 1906. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339372.

Anonymous:1906:VIb

[335] Anonymous. Volume information. Journal of the Royal Statistical Society, 69 (4):823–833, December 1906. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339265.

Martin:1906:ESP

[336] Richard Biddulph Martin and Bart. The electoral “swing of the pendulum”.

Journal of the Royal Statistical Society, 69(4):655–707, December 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339266.

Anonymous:1906:PN

[337] Anonymous. Proceedings on the 20th November, 1906. Journal of the Royal Statistical Society, 69(4):708, December 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339267.

Harris:1906:ERW

[338] William J. Harris and Kenneth A. Lake.

Estimates of the realisable wealth of the United Kingdom based mostly on the estate duty returns. Journal of the Royal Statistical Society, 69(4):709– 745, December 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339268.

Anonymous:1906:CEF

[339] Anonymous. Comparisons of English and foreign agriculture. Journal of the Royal Statistical Society, 69(4):746– 751, December 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339269.

Anonymous:1906:BFC

[340] Anonymous. British and foreign coal in 1905. Journal of the Royal Statistical Society, 69(4):751–752, December 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339270.

Anonymous:1906:PCP

[341] Anonymous. Produce of crops. Preliminary statement showing the estimated total produce and yield per acre of the principal crops in Great Britian in the year 1906, with comparisons for 1905, and the average yield per acre of the ten years 1896–1905. Journal of the Royal Statistical Society, 69(4):753– 754, December 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339271.

Anonymous:1906:DMG

[342] Anonymous. Donation by Mr. G. Udny Yule of works on the theory of probability. Journal of the Royal Statistical Society, 69(4):755–756, December 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339272.

B:1906:BRBh

[343] A. L. B. Book review: Vorlesungen über Mathematische Statistik (Die Lehre von den Statistischen Masszahlen), by Ernst Blaschke. Journal of the Royal Statistical Society, 69(4):757, December 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339273.

B:1906:BRBi

[344] A. L. B. Book review: Instituts Solvay. Travaux de l’Institut de Sociologie. Notes et Memoires. Fascicule 6. Mesure des Capacites Intellectuelle et Energique, by Ch. Henry; E. Waxweiler. Journal of the Royal Statistical Society, 69(4):758–759, December 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339274.

Y:1906:BRBe

[345] G. U. Y. Book review: Mental and Moral Heredity in Royalty, by F. A. Woods. Journal of the Royal Statistical Society, 69(4):759–762, December 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339275.

B:1906:BRBj

[346] A. L. B. Book review: Inter-Temporary Values, or the Distribution of the Produce in Time, by James C. Smith. Journal of the Royal Statistical Society, 69 (4):762–763, December 1906. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339276.

C:1906:BRBd

[347] P. G. C. Book review: The Relations of Rents, Wages, and Profits in Agriculture, and their Bearing on Rural Depopulation, by J. S. Nicholson. Journal of the Royal Statistical Society, 69(4):763– 765, December 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339277.

C:1906:BRBe

[348] P. G. C. Book review: The Return to the Land, by Jules Meline; Justin Mc-Carthy. Journal of the Royal Statistical Society, 69(4):765–768, December 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339278.

H:1906:BRBb

[349] B. L. H. Book review: La Femme dans L’Industrie, by H. Gonnard. Journal of the Royal Statistical Society, 69(4):768– 770, December 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339279.

T:1906:BRB

[350] D. A. T. Book review: The Coal Question: An Inquiry Concerning the Progress of the Nation, and the Probable Exhaustion of our Coal Mines, by W. Stanley Jevons; A. W. Flux. Journal of the Royal Statistical Society, 69(4):770– 772, December 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339280.

B:1906:BRBk

[351] A. L. B. Book review: Imports and Exports at Prices of 1900. Journal of the Royal Statistical Society, 69(4):772– 773, December 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339281.

P:1906:BRBj

[352] L. L. P. Book review: The Wheel of Wealth: Being a Reconstruction of the Science and Art of Political Economy on the Lines of Modern Evolution, by John Beattie Crozier. Journal of the Royal Statistical Society, 69(4):773– 777, December 1906. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339282.

P:1906:BRBk

[353] L. L. P. Book review: Riches and Poverty, by L. G. Chiozza-Money. Journal of the Royal Statistical Society, 69 (4):777–780, December 1906. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339283.

P:1906:BRBl

[354] L. L. P. Book review: The Evolution of Modern Capitalism: A Study of Machine Production, by John A. Hobson. Journal of the Royal Statistical Society, 69(4):780–783, December 1906. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339284.

B:1906:BRBl

[355] E. B. Book review: English Local Government from the Revolution to the Municipal Corporations Act: The Parish and the County, by Sidney Webb; Beatrice Webb. Journal of the Royal Statistical Society, 69(4):783–785, December 1906. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339285.

C:1906:BRBf

[356] M. H. C. Book review: L’Habitation a Bon Marche, by M. Lucien Ferrand; M. Jules Siegfried. Journal of the Royal Statistical Society, 69(4):785–787, December 1906. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339286.

R:1906:BRB

[357] S. R. Book review: Report on the Physical Condition of Fourteen Hundred School Children in the City (of Edinburgh); Together with Some Account of Their Homes and Surroundings (Edinburgh Charity Organisation Society). Journal of the Royal Statistical Society, 69(4):787–791, December 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339287.

Anonymous:1906:ONPc

[358] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 69(4):791–794, December 1906. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339288.

Anonymous:1906:SEAd

[359] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 69(4):795– 803, December 1906. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339289.

Anonymous:1906:QLAd

[360] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 69(4):804–821, December 1906. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339290.

Anonymous:1907:VIa

[361] Anonymous. Volume information. Journal of the Royal Statistical Society, 70 (1):i–vii, March 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339500.

Hooker:1907:CWC

[362] R. H. Hooker. Correlation of the weather and crops. Journal of the Royal Statistical Society, 70(1):1–51, March 1907. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339501.

Yule:1907:SPC

[363] G. Udny Yule. Statistics of production and the census of Production Act (1906). Journal of the Royal Statistical Society, 70(1):52–99, March 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339502.

Craigie:1907:LVG

[364] P. G. Craigie. The Late Viscount Goschen (President 1886–88). Journal of the Royal Statistical Society, 70(1): 100–102, March 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339503.

Edgeworth:1907:RSF

[365] F. Y. Edgeworth. On the representation of statistical frequency by a series. Journal of the Royal Statistical Society, 70(1):102–106, March 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339504.

Sauerbeck:1907:PC

[366] A. Sauerbeck. Prices of commodities in 1906. Journal of the Royal Statistical Society, 70(1):107–121, March 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339505.

Gaster:1907:RFY

[367] Leon Gaster. Roumania’s forty years’ progress, 1866–1906. Journal of the Royal Statistical Society, 70(1):122– 129, March 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339506.

Bailey:1907:C

[368] A. H. Bailey, W. J. H., and K. A. L. Correspondence. Journal of the Royal Statistical Society, 70(1):129–131, March 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339507.

B:1907:BRBa

[369] A. L. B. Book review: Frequency-Curves and Correlation, by W. Palin Elderton; C. Layton; E. Layton. Journal of the Royal Statistical Society, 70(1): 132–134, March 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339508.

B:1907:BRBb

[370] A. L. B. Book review: Employment Pictures from the Census, by M. G. Spencer; H. J. Falk; C. S. Loch. Journal of the Royal Statistical Society, 70(1): 134–135, March 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339509.

N:1907:BRBa

[371] A. N. Book review: Handbuch der Medizinischen Statistik, by Friedrich Prinzing. Journal of the Royal Statistical Society, 70(1):135–136, March 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339510.

C:1907:BRBa

[372] S. J. C. Book review: Cotton: Its Cultivation, Marketing, Manufacture, and the Problems of the Cotton World, by Charles William Burkett; Clarence Hamilton. Journal of the Royal Statistical Society, 70(1):136, March 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339511.

C:1907:BRBb

[373] S. J. C. Book review: From the Cotton Field to the Cotton Mill; a Study of the Industrial Transition in North Carolina, by Holland Thompson. Journal of the Royal Statistical Society, 70(1): 136–138, March 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339512.

P:1907:BRBa

[374] L. L. P. Book review: Des Crises Generales et Periodiques de Surproduction, by Jean Lescure. Journal of the Royal Statistical Society, 70(1):138– 141, March 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339513.

B:1907:BRBc

[375] E. B. Book review: Assistance des Vieillards, Infirmes et Incurables. Commentaire de la Loi Du 14 Juillet, 1905, et des Reglements d’Administration Publique qui en Assurent l’Application. Contenant, en outre, les Instructions Ministérielles et les Circulaires des 16 Avril et 18 Août, 1906, by Adrien Sachet.

Journal of the Royal Statistical Society, 70(1):141–143, March 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339514.

B:1907:BRBd

[376] A. L. B. Book review: Introduction to Economics, by Henry Rogers Seager. Journal of the Royal Statistical Society, 70(1):143–144, March 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339515.

P:1907:BRBb

[377] L. L. P. Book review: On Municipal and National Trading, by Lord Avebury. Journal of the Royal Statistical Society, 70(1):144–146, March 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339516.

P:1907:BRBc

[378] L. L. P. Book review: Municipal Ownership in Great Britain, by Hugo Robert Meyer. Journal of the Royal Statistical Society, 70(1):146–148, March 1907. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339517.

H:1907:BRBa

[379] E. J. H. Book review: Local and Central Government, by Percy Ashley. Journal of the Royal Statistical Society, 70(1): 148–150, March 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339518.

P:1907:BRBd

[380] L. L. P. Book review: Protective and Preferential Import Duties, by A. C. Pigou. Journal of the Royal Statistical Society, 70(1):150–153, March 1907. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339519.

Anonymous:1907:ONPa

[381] Anonymous. Other new publications. Journal of the Royal Statistical Society, 70(1):153–155, March 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339520.

Anonymous:1907:SEAa

[382] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 70(1): 156–164, March 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339521.

Anonymous:1907:QLAa

[383] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 70(1):165–183, March 1907. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339522.

Anonymous:1907:PR

[384] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 70(1): 184–202, March 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339523.

Humphreys:1907:AII

[385] Noel A. Humphreys. The alleged increase of insanity. Journal of the Royal Statistical Society, 70(2):203–241, June 1907. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339661.

Lark:1907:HF

[386] Albert E. Lark. The Herring fishery. Journal of the Royal Statistical Society, 70(2):242–266, June 1907. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339662.

Eversley:1907:DNA

[387] Lord Eversley. The decline in number of agricultural labourers in Great Britain. Journal of the Royal Statistical Society, 70(2):267–319, June 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339663.

Anonymous:1907:RCF

[388] Anonymous. Report of the Council for the financial year ended 31st December, 1906, and for the sessional year ending 18th June, 1907, presented at the seventy-third Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, W.C., London, on the 18th of June, 1907. Journal of the Royal Statistical Society, 70(2):320–329, June 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339664.

Anonymous:1907:PST

[389] Anonymous. Proceedings of the seventythird Annual General Meeting, held on the 18th June, 1907. Journal of the Royal Statistical Society, 70(2):330–331, June 1907. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339665.

Korosy:1907:DJK

[390] Joseph Korosy. Dr. Joseph Korosy — M. G. Olanesco. Journal of the Royal Statistical Society, 70(2):332–333, June 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339666.

Y:1907:BRBa

[391] G. U. Y. Book review: Statistique et Science Sociale: Apercus Generaux. Conferences Données a l’Institut Supérieur de Philosophie de Louvain, by Camille Jacquart. Journal of the Royal Statistical Society, 70(2):334–335, June 1907. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339667.

Y:1907:BRBb

[392] G. U. Y. Book review: Probability, the Foundation of Eugenics. (The Herbert Spencer Lecture.), by Francis Galton. Journal of the Royal Statistical Society, 70(2):335, June 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339668.

C:1907:BRBc

[393] R. F. C. Book review: Annuario Statistico Della Citta Italiane. Journal of the Royal Statistical Society, 70(2):335–336, June 1907. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339669.

Y:1907:BRBc

[394] G. U. Y. Book review: The Inheritance of Ability, by Edgar Schuster. Journal of the Royal Statistical Society, 70(2): 336–337, June 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339670.

N:1907:BRBb

[395] A. N. Book review: De la Frequence des Principales Causes de Deces a Paris Pendant la Seconde Moitie du XIX^{e} Siecle et Notamment Pendant la Periode 1886–1905, by Jacques Bertillon. Journal of the Royal Statistical Society, 70 (2):337–338, June 1907. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339671.

Y:1907:BRBd

[396] G. U. Y. Book review: A First Study of the Statistics of Pulmonary Tuberculosis. (Drapers’ Company Research Memoirs: Studies on National Deterioration, II), by Karl Pearson. Journal of the Royal Statistical Society, 70(2):338, June 1907. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339672.

F:1907:BRBa

[397] O. F. Book review: The Trust Movement in British Industry, by Henry W. Macrosty. Journal of the Royal Statistical Society, 70(2):338–340, June 1907. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339673.

F:1907:BRBb

[398] O. F. Book review: Industrial Combination, by D. H. Macgregor. Journal of the Royal Statistical Society, 70(2):340–344, June 1907. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339674.

C:1907:BRBd

[399] R. F. C. Book review: Le Condizioni Generali Della Classe Operaia in Milano: Salari, Giornate di Lavoro, Reddito, ecc. Journal of the Royal Statistical Society, 70(2):344–345, June 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339675.

H:1907:BRBb

[400] B. L. H. Book review: Child Labor in the United States, by Census Bureau. Journal of the Royal Statistical Society, 70(2):346–347, June 1907. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339676.

B:1907:BRBe

[401] A. L. B. Book review: Life and Labour in Germany, by Gainsborough Commission. Journal of the Royal Statistical Society, 70(2):347–349, June 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339677.

P:1907:BRBe

[402] L. L. P. Book review: Sweated Industry and the Minimum Wage, by Clementina Black; A. G. Gardiner. Journal of the Royal Statistical Society, 70(2):349–351, June 1907. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339678.

P:1907:BRBf

[403] L. L. P. Book review: The Nature of Capital and Income, by Irving Fisher. Journal of the Royal Statistical Society, 70(2):351–353, June 1907. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339679.

B:1907:BRBf

[404] E. B. Book review: Co-Operative Banking: Its Principles and Practice; with a Chapter on Co-Operative Mortgage Credit, by Henry W. Wolff. Journal of the Royal Statistical Society, 70(2): 353–356, June 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339680.

S:1907:BRBa

[405] J. S. Book review: The Gold Supply and Prosperity, by Byron W. Holt. Journal of the Royal Statistical Society, 70(2): 356–358, June 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339681.

B:1907:BRBg

[406] A. L. B. Book review: Die Staats- und Kommunal-Besteuerung in Deutschland, England, den Vereinigten Staaten von Nordamerika und den Englischen Kolonien, by Eugen Lauterbach. Journal of the Royal Statistical Society, 70 (2):358–359, June 1907. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339682.

Y:1907:BRBe

[407] G. U. Y. Book review: Die Weltwirthschaft: Ein Jahr-und Lesebuch: Unter Mitwirkung Zahlreicher Fachleute, by Ernst von Halle. Journal of the Royal Statistical Society, 70(2):359–360, June 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339683.

C:1907:BRBe

[408] P. G. C. Book review: La Russie Agricole Devant la Crise Agraire, by Alexis Yermoloff. Journal of the Royal Statistical Society, 70(2):360–362, June 1907. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339684.

C:1907:BRBf

[409] P. G. C. Book review: L’Argentine au XX^{e} Siecle, by Albert B. Martinez; Maurice Lewandowski. Journal of the Royal Statistical Society, 70(2):363–364, June 1907. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339685.

P:1907:BRBg

[410] L. L. P. Book review: The Industrial Organisation of an Indian Province, by Theodore Morison. Journal of the Royal Statistical Society, 70(2):364–366, June 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339686.

P:1907:BRBh

[411] L. L. P. Book review: India and Imperial Preference, by Roper Lethbridge. Journal of the Royal Statistical Society, 70(2):366–368, June 1907. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339687.

C:1907:BRBg

[412] P. G. C. Book review: The Domesday Inquest, by Adolphus Ballard. Journal of the Royal Statistical Society, 70(2): 368–370, June 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339688.

C:1907:BRBh

[413] P. G. C. Book review: Land Reform, Occupying Ownership, Peasant Proprietary, and Rural Education, by Jesse Collings. Journal of the Royal Statistical Society, 70(2):370–372, June 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339689.

C:1907:BRBi

[414] P. G. C. Book review: The English Peasantry and the Enclosure of Common Fields, by George Slater. Journal of the Royal Statistical Society, 70(2):372–375, June 1907. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339690.

Anonymous:1907:ONPb

[415] Anonymous. Other new publications. Journal of the Royal Statistical Society, 70(2):375–376, June 1907. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339691.

Anonymous:1907:SEAb

[416] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 70(2): 377–386, June 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339692.

Anonymous:1907:QLAb

[417] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 70(2):387–409, June 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339693.

Adams:1907:SCR

[418] W. G. S. Adams. Some considerations relating to the position of the small holding in the United Kingdom. Journal of the Royal Statistical Society, 70(3):411– 448, September 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339616.

Anonymous:1907:CIS

[419] Anonymous. Congress of the International Statistical Institute. Journal of the Royal Statistical Society, 70(3):449– 454, September 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339617.

Dudfield:1907:NMT

[420] Reginald Dudfield. Note on the mortality from tuberculosis from 1851–1905. Journal of the Royal Statistical Society, 70(3):454–470, September 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339618.

Anonymous:1907:MBA

[421] Anonymous. The meeting of the British Association. Discussion on modern methods of treating observations. Journal of the Royal Statistical Society, 70(3):471–476, September 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339619.

Anonymous:1907:ARGa

[422] Anonymous. Agricultural returns of Great Britain, 1907. Journal of the Royal Statistical Society, 70(3):477–478, September 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339620.

N:1907:BRBc

[423] A. N. Book review: Supplement to the Sixty-Fifth Annual Report of the Registrar-General of Births, Deaths and Marriages in England and Wales, 1891– 1900. Journal of the Royal Statistical Society, 70(3):479–481, September 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339621.

R:1907:BRBa

[424] S. R. Book review: Report on the Trade in Imports and Exports at Irish Ports in 1905. Journal of the Royal Statistical Society, 70(3):481–483, September 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339622.

B:1907:BRBh

[425] E. B. Book review: Old Age Pensions. Tables which have Been Prepared in Connection with the Question of Old Age Pensions, with a Preliminary Memorandum. Journal of the Royal Statistical Society, 70(3):483–485, September 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339623.

Y:1907:BRBf

[426] G. U. Y. Book review: Report on a Collection of Statistics as to the Physical Condition of Children Attending the Public Schools of the School Board for Glasgow, by W. Leslie Mackenzie; A.

Foster. Journal of the Royal Statistical Society, 70(3):485–486, September 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339624.

F:1907:BRBc

[427] A. W. F. Book review: Census and Statistics (Canada). Bulletin I. Wage-Earners by Occupations, by S. E. Dawson. Journal of the Royal Statistical Society, 70(3):486–489, September 1907. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339625.

B:1907:BRBi

[428] A. L. B. Book review: Album Graphique de la Statistique Generale de la France. Journal of the Royal Statistical Society, 70(3):489–490, September 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339626.

Y:1907:BRBg

[429] G. U. Y. Book review: Statistiches Handbuch fur das Deutsche Reich. Journal of the Royal Statistical Society, 70 (3):490–491, September 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339627.

Y:1907:BRBh

[430] G. U. Y. Book review: Uebersichten aus der Berliner Statistik fur das Jahr 1905. Journal of the Royal Statistical Society, 70(3):491, September 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339628.

N:1907:BRBd

[431] A. N. Book review: Modern Social Conditions. A Statistical Study of Birth, Marriage, Divorce, Death, Disease, Suicide, Immigration, with Special Reference to the United States, by Wm. B. Bailey. Journal of the Royal Statistical Society, 70(3):491–492, September 1907. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339629.

Y:1907:BRBi

[432] G. U. Y. Book review: Heredity and Selection in Sociology, by George Chatterton-Hill. Journal of the Royal Statistical Society, 70(3):492– 493, September 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339630.

Y:1907:BRBj

[433] G. U. Y. Book review: Mendelism, by R. C. Punnett. Journal of the Royal Statistical Society, 70(3):493–494, September 1907. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339631.

Y:1907:BRBk

[434] G. U. Y. Book review: Studien zur Bevolkerungsbewegung in Deutschland, by Paul Mombert. Journal of the Royal Statistical Society, 70(3):494, September 1907. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339632.

N:1907:BRBe

[435] A. N. Book review: La Mortalite Infantile dans les Flanders: Etude de Demographie Belge, by Camille Jacquart. Journal of the Royal Statistical Society, 70(3):494–495, September 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339633.

B:1907:BRBj

[436] E. B. Book review: The Town Child, by Reginald A. Bray. Journal of the Royal Statistical Society, 70(3):495– 497, September 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339634.

H:1907:BRBc

[437] E. J. H. Book review: The Housing Problem in England, 1907, by E. R. Dewsnup. Journal of the Royal Statistical Society, 70(3):497, September 1907. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339635.

H:1907:BRBd

[438] E. J. H. Book review: Housing, 1907, by P. Alden; E. E. Hayward. Journal of the Royal Statistical Society, 70(3):498– 499, September 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339636.

H:1907:BRBe

[439] E. J. H. Book review: Le Surpeuplement et les Habitations a Bon Marche, 1907, by H. Turot; H. Bellamy. Journal of the Royal Statistical Society, 70(3): 499, September 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339637.

B:1907:BRBk

[440] E. B. Book review: Le Socialisme Municipal en Angleterre et ses Resultats Financiers, by Raymond Boverat. Journal of the Royal Statistical Society, 70(3): 499–501, September 1907. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339638.

F:1907:BRBd

[441] A. W. F. Book review: The Dangers of Municipal Trading, by Robert P. Porter. Journal of the Royal Statistical Society, 70(3):501–504, September 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339639.

P:1907:BRBi

[442] L. L. P. Book review: Political Economy, by Charles S. Devas. Journal of the Royal Statistical Society, 70(3):504– 505, September 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339640.

P:1907:BRBj

[443] L. L. P. Book review: La Science Economique, by Yves Guyot. Journal of the Royal Statistical Society, 70(3):505– 507, September 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339641.

P:1907:BRBk

[444] L. L. P. Book review: Economie Politique, by Felix Martel; M. Grigaut. Journal of the Royal Statistical Society, 70 (3):507–508, September 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339642.

B:1907:BRBl

[445] A. L. B. Book review: Le Salaire des Ouvriers des Mines de Charbon en France, by Francois Simiand. Journal of the Royal Statistical Society, 70(3):508– 509, September 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339643.

P:1907:BRBl

[446] L. L. P. Book review: The Strength of Nations: An Argument from History, by J. W. Welsford. Journal of the Royal Statistical Society, 70(3):509– 511, September 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339644.

C:1907:BRBj

[447] P. G. C. Book review: The Small Holdings of England: A Survey of the Various Existing Systems, by L. Jebb. Journal of the Royal Statistical Society, 70(3):512–515, September 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339645.

P:1907:BRBm

[448] L. L. P. Book review: A History of William Paterson and the Darien Company, by James Samuel Barbour. Journal of the Royal Statistical Society, 70 (3):515–517, September 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339646.

Y:1907:BRBl

[449] G. U. Y. Book review: Sociological Papers. Journal of the Royal Statistical Society, 70(3):517–518, September 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339647.

Anonymous:1907:ONPc

[450] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 70(3):519–520, September 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339648.

Anonymous:1907:SEAc

[451] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 70(3):521– 529, September 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339649.

Anonymous:1907:QLAc

[452] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 70(3):530–551, September 1907. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339650.

Anonymous:1907:VIb

[453] Anonymous. Volume information. Journal of the Royal Statistical Society, 70 (4):729–743, December 1907. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339559.

Dilke:1907:PAR

[454] Charles W. Dilke. The Presidential address of the Right Hon. Sir Charles W. Dilke, Bart., M.P., for the session 1907– 08. Delivered to the Royal Statistical Society, 19th November, 1907. Journal of the Royal Statistical Society, 70(4):553– 582, December 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339560.

Anonymous:1907:PN

[455] Anonymous. Proceedings on the 19th November, 1907. Journal of the Royal Statistical Society, 70(4):583– 586, December 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339561.

Thompson:1907:IRA

[456] Robert J. Thompson. An inquiry into the rent of agricultural land in England and Wales during the Nineteenth Century. Journal of the Royal Statistical Society, 70(4):587–624, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339562.

Welton:1907:MBR

[457] Thomas A. Welton. Memorandum as to birth-rates and marriage-rates in England and Wales. Journal of the Royal Statistical Society, 70(4):625– 629, December 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339563.

Ballard:1907:ABO

[458] Adolphus Ballard. The assize of bread at Oxford, 1794–1820. Journal of the Royal Statistical Society, 70(4):629– 637, December 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339564.

Moore:1907:DLW

[459] Henry L. Moore. The differential law of wages. Journal of the Royal Statistical Society, 70(4):638–651, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339565.

Anonymous:1907:ARGb

[460] Anonymous. Agricultural returns of Great Britain, 1907. Journal of the Royal Statistical Society, 70(4):651– 653, December 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339566.

Anonymous:1907:ICT

[461] Anonymous. International Congress on tuberculosis. Journal of the Royal Statistical Society, 70(4):654, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339567.

Y:1907:BRBm

[462] G. U. Y. Book review: On Further Methods of Determining Correlation, by Karl Pearson. Journal of the Royal Statistical Society, 70(4):655–656, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339568.

C:1907:BRBk

[463] C. E. C. Book review: Statistics of Women at Work; Based on Unpublished Information Derived from the Schedules of the Twelfth Census, 1900. Journal of the Royal Statistical Society, 70(4):656– 659, December 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339569.

Y:1907:BRBn

[464] G. U. Y. Book review: Wage-Earners’ Budgets: A Study of Standards and Cost of Living in New York City, by Louise Bolard More. Journal of the Royal Statistical Society, 70(4):659–661, December 1907. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339570.

S:1907:BRBb

[465] C. P. S. Book review: Die Statistischen Unterlagen der Oesterreichischen Wahlreform, by H. Rauchberg. Journal of the Royal Statistical Society, 70(4): 661–662, December 1907. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339571.

N:1907:BRBf

[466] A. N. Book review: The Classification of Disease and Causes of Death, from the Standpoint of the Statistician, by G. H. Knibbs. Journal of the Royal Statistical Society, 70(4):662–663, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339572.

D:1907:BRB

[467] R. D. Book review: Unification of Australasian Statistical Methods, and Co-Ordination of the Work of the Commonwealth and State Bureaux. Journal of the Royal Statistical Society, 70(4):663– 664, December 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339573.

C:1907:BRBl

[468] P. G. C. Book review: Meat Supply and Surplus, with Consideration of Consumption and Exports, by George K. Holmes. Journal of the Royal Statistical Society, 70(4):664–667, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339574.

Y:1907:BRBo

[469] G. U. Y. Book review: A First Study of the Statistics of Insanity and the Inheritance of the Insane Diathesis, by David Heron. Journal of the Royal Statistical Society, 70(4):667, December 1907. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339575.

P:1907:BRBn

[470] L. L. P. Book review: Board of Agriculture and Fisheries. Report on the Production of Wool in Great Britain in 1905 and 1906. Journal of the Royal Statistical Society, 70(4):668–670, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339576.

N:1907:BRBg

[471] A. N. Book review: Physical Welfare of School Children. An Examination of the Home Conditions of 1,400 New York School Children Found, by School Physicians to Have Physical Defects. by New York Committee on Physical Welfare of School Children. Journal of the Royal Statistical Society, 70 (4):670–671, December 1907. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339577.

F:1907:BRBe

[472] A. W. F. Book review: The Tariff Commission. Vol. 3. Report of the Agricultural Committee. Journal of the Royal Statistical Society, 70(4):671– 674, December 1907. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339578.

P:1907:BRBo

[473] L. L. P. Book review: Questions Ouvrieres et Industrielles en France sous la Troisieme Republique, by E. Levasseur. Journal of the Royal Statistical Society, 70(4):674–678, December 1907. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339579.

P:1907:BRBp

[474] L. L. P. Book review: La Monnaie, by A. de Foville. Journal of the Royal Statistical Society, 70(4):678–680, December 1907. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339580.

P:1907:BRBq

[475] L. L. P. Book review: L’Impot sur le Revenu en Allemagne, by Gaspard Wampach. Journal of the Royal Statistical Society, 70(4):680–682, December 1907. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339581.

B:1907:BRBm

[476] P. E. B. Book review: Untersuchung der Wohnungen der Minder Bemittelten Klassen in Bremen: im Auftrage der Deputation wegen Wohnungsinspektion Veranstaltet im April 1904 und Bearbeitet vom Bremischen Statistischen Amt. Journal of the Royal Statistical Society, 70(4):683–686, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339582.

F:1907:BRBf

[477] A. W. F. Book review: Practical Problems in Banking and Currency, by Walter Henry Hull. Journal of the Royal Statistical Society, 70(4):686–688, December 1907. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339583.

C:1907:BRBm

[478] R. F. C. Book review: Le Crisi Economiche, by Di Camillo Supino. Journal of the Royal Statistical Society, 70 (4):688–690, December 1907. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339584.

R:1907:BRBb

[479] S. R. Book review: Bevolkerungsstatistik Schwedens (1750–1900), by Gustav Sundbarg. Journal of the Royal Statistical Society, 70(4):690–692, December 1907. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339585.

Anonymous:1907:C

[480] Anonymous. Correspondence. Journal of the Royal Statistical Society, 70(4): 692–693, December 1907. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339586.

Anonymous:1907:ONPd

[481] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 70(4):693–695, December 1907. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339587.

Anonymous:1907:SEAd

[482] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 70(4):696– 705, December 1907. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339588.

Anonymous:1907:QLAd

[483] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 70(4):706–727, December 1907. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339589.

Anonymous:1908:VIa

[484] Anonymous. Volume information. Journal of the Royal Statistical Society, 71 (1):i–vii, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339242.

Dudfield:1908:SUF

[485] Reginald Dudfield. Some unconsidered factors affecting the birth-rate. Journal of the Royal Statistical Society, 71 (1):1–64, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339243.

Mallet:1908:MEC

[486] Bernard Mallet. A method of estimating capital wealth from the estate duty statistics. Journal of the Royal Statistical Society, 71(1):65–101, March 1908. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339244.

Edwards:1908:RTG

[487] Charles Lewis Edwards. Railways and the trade of Great Britain. Journal of the Royal Statistical Society, 71(1): 102–131, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339245.

Sauerbeck:1908:PC

[488] A. Sauerbeck. Prices of commodities in 1907. Journal of the Royal Statistical Society, 71(1):132–146, March 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339246.

Cook:1908:EAD

[489] J. Basil Cook. An examination of the amount of in-door pauperism in three metropolitan boroughs, and of the causes which to led to this pauperism. Journal of the Royal Statistical Society, 71(1):147–174, March 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339247.

Hutchins:1908:NMY

[490] B. L. Hutchins. Note on the mortality of Young children. Journal of the Royal Statistical Society, 71(1):174– 178, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339248.

Anonymous:1908:BPC

[491] Anonymous. Bibliography of the prices of cereals. Presented by the Royal Statistical Society to the International Statistical Institute at the Copenhagen Congress, August, 1907. Journal of the Royal Statistical Society, 71(1):178– 206, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339249.

H:1908:BRBa

[492] E. J. H. Book review: Report of an Enquiry, by the Board of Trade into Working Class Rents, Housing and Retail Prices, together with the Standard Rates of Wages Prevailing in Certain Occupations in the Principal Industrial Towns of the United Kingdom. Journal of the Royal Statistical Society, 71(1): 207–213, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339250.

B:1908:BRBa

[493] A. L. B. Book review: Amalgamated Society of Railway Servants. Report Giving Statement of the Results of the Census of Wages, Hours of Labour, &c., of Men Employed on the Railways of the United Kingdom. Journal of the Royal Statistical Society, 71(1):213–214, March 1908. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339251.

K:1908:BRB

[494] J. M. K. Book review: West Ham: A Study in Social and Industrial Problems, by Edward G. Howarth. Journal of the Royal Statistical Society, 71(1): 215–217, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339252.

H:1908:BRBb

[495] B. L. H. Book review: Cambridge: A Brief Study in Social Questions, by Eglantyne Jebb. Journal of the Royal Statistical Society, 71(1):217– 218, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339253.

C:1908:BRBa

[496] C. E. C. Book review: L’Ouvriere en France, by Caroline Milhaud. Journal of the Royal Statistical Society, 71(1): 218–220, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339254.

H:1908:BRBc

[497] B. L. H. Book review: Die Deutsche Frau um die Jahrhundertwende, by Elis Gnauck-Kuhne. Journal of the Royal Statistical Society, 71(1):221– 222, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339255.

C:1908:BRBb

[498] G. G. C. Book review: Sozialstatistik, by Gottlieb Schnapper-Arndt. Journal of the Royal Statistical Society, 71(1): 222–223, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339256.

P:1908:BRBa

[499] L. L. P. Book review: Municipal and Private Operation of Public Utilities. Journal of the Royal Statistical Society, 71(1):223–229, March 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339257.

Fl:1908:BRBa

[500] A. W. Fl. Book review: Essentials of Economic Theory, by John Bates Clark. Journal of the Royal Statistical Society, 71(1):229–230, March 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339258.

Fl:1908:BRBb

[501] A. W. Fl. Book review: The Rate of Interest, by Irving Fisher. Journal of the Royal Statistical Society, 71(1): 230–231, March 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339259.

C:1908:BRBc

[502] G. G. C. Book review: Anglo-Chinese Commerce and Diplomacy, by A. J. Sargent. Journal of the Royal Statistical Society, 71(1):231–233, March 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339260.

Anonymous:1908:ONPa

[503] Anonymous. Other new publications. Journal of the Royal Statistical Society, 71(1):233–236, March 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339261.

Anonymous:1908:SEAa

[504] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 71(1): 237–246, March 1908. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339262.

Anonymous:1908:QLAa

[505] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 71(1):247–271, March 1908. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339263.

Anonymous:1908:PR

[506] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 71(1): 272–289, March 1908. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339264.

Baines:1908:PIL

[507] J. Athelstane Baines. The peradventures of an Indian life-table. Journal of the Royal Statistical Society, 71(2):291–318, June 1908. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339457.

Rosenbaum:1908:FTU

[508] S. Rosenbaum. Food taxation in the United Kingdom, France, Germany, and the United States. Journal of the Royal Statistical Society, 71(2):319–365, June 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339458.

Anonymous:1908:RCF

[509] Anonymous. Report of the Council for the financial year ended 31st December, 1907, and for the sessional year ending 16th June, 1908, presented at the seventy-fourth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, W.C., London, on the 16th of June, 1908. Journal of the Royal Statistical Society, 71(2):366–377, June 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339459.

Anonymous:1908:PSF

[510] Anonymous. Proceedings of the seventyfourth Annual General Meeting, held on 16th June, 1908. Journal of the Royal Statistical Society, 71(2):378–380, June 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339460.

Edgeworth:1908:PEFa

[511] F. Y. Edgeworth. On the probable errors of frequency-constants. Journal of the Royal Statistical Society, 71(2):381–397, June 1908. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339461.

Godfrey:1908:SAD

[512] Ernest H. Godfrey. Settlement and agricultural development of the North-West provinces of Canada. Journal of the Royal Statistical Society, 71(2):397–404, June 1908. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339462.

Y:1908:BRBa

[513] G. U. Y. Book review: Statistical Reorganisation. Report of a Committee Appointed, by the Secretary of Commerce and Labor to Inquire into the Statistical Work of the Department. Journal of the Royal Statistical Society, 71(2):405–406, June 1908. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339463.

Y:1908:BRBb

[514] G. U. Y. Book review: Statistique Mathématique, by H. Laurent. Journal of the Royal Statistical Society, 71(2): 406–407, June 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339464.

C:1908:BRBd

[515] P. G. C. Book review: Cereal Production of Europe by U.S. Department of Agriculture, Bureau of Statistics; European Grain Trade by U.S. Department of Agriculture, Bureau of Statistics; Russia’s Wheat Trade, by U.S. Department of Agriculture, Bureau of Statistics. Journal of the Royal Statistical Society, 71(2):407–409, June 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339465.

Fl:1908:BRBc

[516] A. W. Fl. Book review: Money and Credit Instruments in their Relation to General Prices, by E. W. Kemmerer. Journal of the Royal Statistical Society, 71(2):409–410, June 1908. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339466.

C:1908:BRBe

[517] R. F. C. Book review: Wirtschaftskrisen und Ueberkapitalisation, by Mentor Bouniatian. Journal of the Royal Statistical Society, 71(2):410–413, June 1908. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339467.

Fl:1908:BRBd

[518] A. W. Fl. Book review: L’Industrie Américaine, by Achille Viallate. Journal of the Royal Statistical Society, 71(2): 413–414, June 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339468.

Fl:1908:BRBe

[519] A. W. Fl. Book review: American Business Enterprise. A Study in Industrial Organisation, by Douglas Knoop. Journal of the Royal Statistical Society, 71 (2):414–415, June 1908. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339469.

P:1908:BRBb

[520] L. L. P. Book review: The Economic History of the United States, by Ernest Ludlow Bogart. Journal of the Royal Statistical Society, 71(2):415–416, June 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339470.

B:1908:BRBb

[521] E. B. Book review: Les Traités Ouvriers: Accords Internationaux de Prévoyance et de Travail. Textes Officiels. Commentaire et Historique, by Albert Métin. Journal of the Royal Statistical Society, 71(2):416–417, June 1908. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339471.

H:1908:BRBd

[522] E. J. H. Book review: The Sanitary Evolution of London, by Henry Jephson. Journal of the Royal Statistical Society, 71(2):417–418, June 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339472.

A:1908:BRB

[523] P. A. Book review: English Local Government from the Revolution to the Municipal Corporations Act: The Manor and the Borough, by Sidney Webb; Beatrice Webb. Journal of the Royal Statistical Society, 71(2):418–420, June 1908. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339473.

Anonymous:1908:ONPb

[524] Anonymous. Other new publications. Journal of the Royal Statistical Society, 71(2):420–423, June 1908. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339474.

Anonymous:1908:SEAb

[525] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 71(2): 424–432, June 1908. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339475.

Anonymous:1908:QLAb

[526] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 71(2):433–458, June 1908. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339476.

Bowley:1908:IOS

[527] Arthur L. Bowley. The improvement of official statistics. Journal of the Royal Statistical Society, 71(3):459–495, September 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339291.

Anonymous:1908:RCC

[528] Anonymous. The report of the census committee. Journal of the Royal Statistical Society, 71(3):496–498, September 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339292.

Edgeworth:1908:PEFb

[529] F. Y. Edgeworth. On the probable errors of frequency-constants (contd.). Journal of the Royal Statistical Society, 71(3): 499–512, September 1908. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339293.

Collet:1908:SSW

[530] Clara E. Collet. The social status of women occupiers. Journal of the Royal Statistical Society, 71(3):513– 515, September 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339294.

Layton:1908:CWD

[531] W. T. Layton. Changes in the wages of domestic servants during fifty years. Journal of the Royal Statistical Society, 71(3):515–524, September 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339295.

Anonymous:1908:ARGa

[532] Anonymous. Agricultural returns of Great Britain, 1908. Journal of the Royal Statistical Society, 71(3):525–526, September 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339296.

H:1908:BRBe

[533] E. J. H. Book review: Report of an Enquiry, by the Board of Trade into Working-Class Rents, Housing and Retail Prices, Together with the Rates of Wages in Certain Occupations in the Principal Industrial Towns of the German Empire. Journal of the Royal Statistical Society, 71(3):527–531, September 1908. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339297.

Y:1908:BRBc

[534] G. U. Y. Book review: Ein Mittelbürgerliches Budget über Einen Zehnjährigen Zeitraum, nebst Anhang “Die Verteuerung der Lebenshaltung im Lichte des Massenkonsums.”, by Henriette Fürth. Journal of the Royal Statistical Society, 71(3):531–532, September 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339298.

R:1908:BRB

[535] J. A. R. Book review: Statistique Internationale du Mouvement de la Population, d’après les Registres d’Etat Civil. (Statistique Générale de la France, 1907.). Journal of the Royal Statistical Society, 71(3):532–535, September 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339299.

Y:1908:BRBd

[536] G. U. Y. Book review: Statistique Annuelle du Mouvement de la Population de la France d’après les Registres d’état Civil au Cours de l’année 1907. Journal of the Royal Statistical Society, 71(3): 536, September 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339300.

Y:1908:BRBe

[537] G. U. Y. Book review: Beiträge zum Problem der Volksichte, by Richard Tronnier. Journal of the Royal Statistical Society, 71(3):536, September 1908. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339301.

Y:1908:BRBf

[538] G. U. Y. Book review: Official Year-Book of the Commonwealth of Australia, Containing Authoritative Statistics for the Period 1901–07, and Corrected Statistics for the Period 1788– 1900, by G. H. Knibbs. Journal of the Royal Statistical Society, 71(3):537– 538, September 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339302.

Y:1908:BRBg

[539] G. U. Y. Book review: The Brazilian Year-Book, Issued Under the Patronage of the Brazilian Government, by J. P. Wileman. Journal of the Royal Statistical Society, 71(3):538, September 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339303.

H:1908:BRBf

[540] N. A. H. Book review: United Kingdom. Report of the Royal Commission on the Care and Control of the Feeble-Minded. Journal of the Royal Statistical Society, 71(3):539–543, September 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339304.

C:1908:BRBf

[541] P. G. C. Book review: Report of the Select Committee on the Hop Industry. Journal of the Royal Statistical Society, 71(3):543–547, September 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339305.

C:1908:BRBg

[542] W. H. C. Book review: Report, by the Committee on a National Guarantee for the War Risks of Shipping to the Lords Commissioners of H. M. Treasury, 1908. Journal of the Royal Statistical Society, 71(3):547–549, September 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339306.

C:1908:BRBh

[543] R. I. C. Book review: Geschichte der Handelskrisen in England, by Mentor Bouniatian. Journal of the Royal Statistical Society, 71(3):549–551, September 1908. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339307.

H:1908:BRBg

[544] E. J. H. Book review: La Vie des Grandes Capitales: Études Comparatives sur Londres, Paris, Berlin, by Gaston Cadoux. Journal of the Royal Statistical Society, 71(3):551–552, September 1908. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339308.

S:1908:BRB

[545] J. C. S. Book review: Essays in Municipal Administration, by John A. Fairlie. Journal of the Royal Statistical Society, 71(3):552–554, September 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339309.

P:1908:BRBc

[546] L. L. P. Book review: La Réforme Fiscale par l’Impôt sur le Revenu, by A.- E. Gauthier; L’Impôt sur le Revenu en Prusse, by René Bomboy; L’Impôt sur le Revenu en Autriche, by Victor Marcé. Journal of the Royal Statistical Society, 71(3):554–559, September 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339310.

F:1908:BRBa

[547] A. F. Book review: Gold, Prices, and Wages under the Greenback Standard, by Wesley C. Mitchell. Journal of the Royal Statistical Society, 71(3):559– 560, September 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339311.

F:1908:BRBb

[548] A. F. Book review: The Fall in Consols and Other Investments since 1897, by A. H. Gibson; Bank Rate, the Banker’s Vade Mecum, by A. H. Gibson. Journal of the Royal Statistical Society, 71(3): 560–561, September 1908. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339312.

F:1908:BRBc

[549] A. F. Book review: Die Wertzuwachssteuer, by Jur. Boldt. Journal of the Royal Statistical Society, 71(3):561– 563, September 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339313.

Y:1908:BRBh

[550] G. U. Y. Book review: Memoir on the Pigmentation Survey of Scotland, by John Gray. Journal of the Royal Statistical Society, 71(3):563, September 1908. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339314.

Y:1908:BRBi

[551] G. U. Y. Book review: A Second Study of the Statistics of Pulmonary Tuberculosis: Marital Infection, by E. G. Pope; Karl Pearson; Ethel M. Elderton. Journal of the Royal Statistical Society, 71 (3):564, September 1908. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339315.

Y:1908:BRBj

[552] G. U. Y. Book review: Ricerche Demografiche Fondate sul Calcolo delle Variazioni, by Luigi Perozzo. Journal of the Royal Statistical Society, 71(3): 564, September 1908. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339316.

C:1908:BRBi

[553] W. H. C. Book review: The Trade and Administration of the Chinese Empire, by Hosea Ballou Morse. Journal of the Royal Statistical Society, 71(3):564– 565, September 1908. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339317.

P:1908:BRBd

[554] L. L. P. Book review: The Early History of the Levant Company, by M. Epstein. Journal of the Royal Statistical Society, 71(3):565–566, September 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339318.

Anonymous:1908:ONPc

[555] Anonymous. Other new publications. Journal of the Royal Statistical Society, 71(3):566–568, September 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339319.

Anonymous:1908:SEAc

[556] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 71(3):569– 577, September 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339320.

Anonymous:1908:QLAc

[557] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 71(3):578–599, September 1908. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339321.

Anonymous:1908:VIb

[558] Anonymous. Volume information. Journal of the Royal Statistical Society, 71 (4):771–788, December 1908. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339373.

Brabrook:1908:SI

[559] Edward Brabrook. Social insurances.

Journal of the Royal Statistical Society, 71(4):601–612, December 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339374. See discussion

[560].

Anonymous:1908:DSE

[560] Anonymous. Discussion on Sir Edward Brabrook’s paper. Journal of the Royal Statistical Society, 71(4):612– 618, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2339375. See [559].

Chapman:1908:AMR

[561] Samuel Chapman. American methods of railway accounting. Journal of the Royal Statistical Society, 71(4):619– 648, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339376.

Anonymous:1908:DIS

[562] Anonymous. Dr. von Inama-Sternegg. Journal of the Royal Statistical Society, 71(4):649–651, December 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339377.

Edgeworth:1908:PEFc

[563] F. Y. Edgeworth. On the probable errors of frequency-constants (contd.). Journal of the Royal Statistical Society, 71(4): 651–678, December 1908. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339378.

Drake:1908:SIL

[564] Mrs. Drake. A study of infant life in Westminster. Journal of the Royal Statistical Society, 71(4):678–686, December 1908. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339379.

Edser:1908:NNP

[565] T. Edser. Note on the number of plaice at each length, in certain samples from the southern part of the North Sea, 1906. Journal of the Royal Statistical Society, 71(4):686–690, December 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339380.

Dudfield:1908:NCT

[566] Reginald Dudfield. A note on certain tables in Part II of the Supplement to the Sixty-Fifth Annual Report of the Registrar-General. Journal of the Royal Statistical Society, 71(4):690– 691, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339381.

Anonymous:1908:ARGb

[567] Anonymous. Agricultural returns of Great Britain, 1908. Journal of the Royal Statistical Society, 71(4):692– 694, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339382.

D:1908:BRBa

[568] R. D. Book review: Supplement to the Sixty-Fifth Annual Report of the Registrar-General of England and Wales. Journal of the Royal Statistical Society, 71(4):695–700, December 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339383.

D:1908:BRBb

[569] R. D. Book review: La Mortalita Infantile in Milano: Risultati di un’ Inchiesta sui Nati nel 1903 in Rapporto ai Modi di Allatamento, e alle Condizioni Economiche dei Genitori. Journal of the Royal Statistical Society, 71(4):700– 703, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339384.

W:1908:BRBa

[570] A. D. W. Book review: Kindersterblichkeit Besonders in den Niederlanden, by R. H. Saltet; Ph. Falkenburg. Journal of the Royal Statistical Society, 71(4):704, December 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339385.

D:1908:BRBc

[571] R. D. Book review: Diseases of Occupation from the Legislative, Social and Medical Points of View, by Thomas Oliver. Journal of the Royal Statistical Society, 71(4):704–705, December 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339386.

P:1908:BRBe

[572] L. L. P. Book review: Report to the Secretary of State for the Home Department on the Wages Board and Industrial Conciliation and Arbitration Acts of Australia and New Zealand, by Ernest Aves; Report from the Select Committee on Home Work Together with the Proceedings. Journal of the Royal Statistical Society, 71(4):705–708, December 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339387.

B:1908:BRBc

[573] A. L. B. Book review: The Central (Unemployed) Body for London. 1. Preliminary Report, 2. Second Report, 3. Special Committee on Unskilled Labour. Report and Minutes of Evidence. Journal of the Royal Statistical Society, 71(4):709– 711, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339388.

B:1908:BRBd

[574] A. L. B. Book review: Statistik der Streiks und Aussperrungen im In- und Auslande, by Maximilian Meyer. Journal of the Royal Statistical Society, 71 (4):711–712, December 1908. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339389.

B:1908:BRBe

[575] E. B. Book review: La Participation aux Benefices en Allemagne, en Autriche et en Suisse, by Victor Bohmert; Albert Trombert. Journal of the Royal Statistical Society, 71(4):712–713, December 1908. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339390.

M:1908:BRB

[576] B. M. Book review: La Richesse de la France. Fortunes et Revenus Prives, by A. de Lavergne; Paul Henry. Journal of the Royal Statistical Society, 71(4):713– 714, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339391.

P:1908:BRBf

[577] L. L. P. Book review: La Progression des Budgets en France du Treizième Siècle à nos Jours: Ses Causes et ses Remèdes, by Joseph Jacob. Journal of the Royal Statistical Society, 71(4):714– 715, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339392.

C:1908:BRBj

[578] P. G. C. Book review: Hungary: A Sketch of the Country, Its People, and Its Conditions, by Julius de Vargha. Journal of the Royal Statistical Society, 71(4):715–716, December 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339393.

F:1908:BRBd

[579] A. F. Book review: The King’s Revenue, by W. M. J. Williams. Journal of the Royal Statistical Society, 71(4):716– 717, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339394.

P:1908:BRBg

[580] L. L. P. Book review: Trade and Tariffs, by John M. Robertson. Journal of the Royal Statistical Society, 71(4):717– 719, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339395.

F:1908:BRBe

[581] A. F. Book review: Die Volkswirthschaftliche Bedeutung der Technischen Entwicklung der Deutschen Zuckerindustrie, by Th. Schuchart. Journal of the Royal Statistical Society, 71(4): 719–721, December 1908. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339396.

C:1908:BRBk

[582] S. J. C. Book review: Der Termin-Handel in Nordamerikanischer Baumwolle, by Carl von Kuhlmann. Journal of the Royal Statistical Society, 71(4):721– 722, December 1908. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339397.

C:1908:BRBl

[583] P. G. C. Book review: A History of the English Agricultural Labourer, by W. Hasbach. Journal of the Royal Statistical Society, 71(4):722–725, December 1908. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339398.

Y:1908:BRBk

[584] G. U. Y. Book review: Zur Frage einer Alkoholconsumstatistik, by Alfred Kubatz. Journal of the Royal Statistical Society, 71(4):725–726, December 1908. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339399.

W:1908:BRBb

[585] A. D. W. Book review: Kraft: Okonomische, Technische und Kulturgeschichtliche Studien über die Machtentfaltung der Staaten, by E. Reyer. Journal of the Royal Statistical Society, 71(4):726–727, December 1908. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339400.

Y:1908:BRBl

[586] G. U. Y. Book review: Adolphe Quetelet as Statistician, by Frank H. Hawkins; Studies in History, Economics, and Public Law, by Faculty of Political Science of Columbia University. Journal of the Royal Statistical Society, 71(4): 728, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339401.

F:1908:BRBf

[587] A. F. Book review: The Financial Status of the Professor in America and in Germany. The Carnegie Foundation for the Advancement of Teaching. Journal of the Royal Statistical Society, 71(4):728– 729, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339402.

P:1908:BRBh

[588] L. L. P. Book review: Collectivism:

A Study of Some of the Leading Social Questions of the Day, by Paul Leroy Beaulieu; Arthur Clay. Journal of the Royal Statistical Society, 71(4):729– 731, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339403.

F:1908:BRBg

[589] A. F. Book review: Value and Distribution; A Critical and Constructive Study, by H. J. Davenport. Journal of the Royal Statistical Society, 71(4):731– 732, December 1908. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339404.

P:1908:BRBi

[590] L. L. P. Book review: The Elements of Economics, by Charles Jesse Bullock. Journal of the Royal Statistical Society, 71(4):732–733, December 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339405.

Anonymous:1908:ONPd

[591] Anonymous. Other new publications. Journal of the Royal Statistical Society, 71(4):734–736, December 1908. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339406.

Anonymous:1908:SEAd

[592] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 71(4):737– 746, December 1908. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339407.

Anonymous:1908:QLAd

[593] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 71(4):747–770, December 1908. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339408.

Anonymous:1909:VIa

[594] Anonymous. Volume information. Journal of the Royal Statistical Society, 72 (1):i–vii, March 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339942.

Braun:1909:CCR

[595] Percy E. Braun. The cost, conditions, and results of hospital relief in London. Journal of the Royal Statistical Society, 72(1):1–39, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339943.

Somerville:1909:FSE

[596] William Somerville. Forestry in some of its economic aspects. Journal of the Royal Statistical Society, 72(1):40– 63, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339944.

B:1909:AWF

[597] A. E. B. Arthur Wilson Fox, C.B. Journal of the Royal Statistical Society, 72 (1):64–66, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339945.

C:1909:CCD

[598] P. G. C. and W. F. Willcox. Colonel Carroll D. Wright. Journal of the Royal Statistical Society, 72(1):66–67, March 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339946.

Sauerbeck:1909:PC

[599] A. Sauerbeck. Prices of commodities in 1908. Journal of the Royal Statistical Society, 72(1):68–80, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339947.

Edgeworth:1909:APE

[600] F. Y. Edgeworth. Addendum on “probable errors of frequency-constants”. Journal of the Royal Statistical Society, 72 (1):81–90, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339948.

Wood:1909:RWS

[601] George H. Wood. Real wages and the standard of comfort since 1850. Journal of the Royal Statistical Society, 72(1): 91–103, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339949.

E:1909:BRB

[602] F. Y. E. Book review: Principii di Statistica Metodologica, by Rodolfo Benini. Journal of the Royal Statistical Society, 72(1):104–110, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339950.

B:1909:BRBa

[603] A. L. B. Book review: Die Statistischen Mittelwerte. Eine Methodologische Untersuchung, by Franz Žižek. Journal of the Royal Statistical Society, 72(1): 110–112, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339951.

B:1909:BRBb

[604] A. L. B. Book review: Théorie du Salaire et du Travail Salarié, by Christian Cornélissen. Journal of the Royal Statistical Society, 72(1):112– 113, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339952.

Y:1909:BRBa

[605] G. U. Y. Book review: Leçons Élémentaires sur le Calcul des Probabilités, by R. de Montessus. Journal of the Royal Statistical Society, 72(1): 113–114, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339953.

H:1909:BRBa

[606] E. J. H. Book review: Report of an Inquiry, by the Board of Trade into Working Class Rents, Housing, and Retail Prices, Together with the Rates of Wages in Certain Occupations in the Principal Industrial Towns of France. Journal of the Royal Statistical Society, 72(1):114–118, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339954.

B:1909:BRBc

[607] E. B. Book review: Royal Commission on the Poor Laws and Relief of Distress. Report Presented to both Houses of Parliament, by Command of His Majesty. Journal of the Royal Statistical Society, 72(1):118–122, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339955.

Anonymous:1909:RCCa

[608] Anonymous. Royal Commission on Coast Erosion and Afforestation. Second report (on afforestation). Journal of the Royal Statistical Society, 72(1): 123–126, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339956.

B:1909:BRBd

[609] E. B. Book review: L’Assistance Obligatoire aux Vieillards, aux Infirmes, et aux Incurables. Commentaire de la loi du 14 Juillet, 1905, by Edouard Campagnole; M. Bienvenu-Martin. Journal of the Royal Statistical Society, 72(1): 126–127, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339957.

B:1909:BRBe

[610] E. B. Book review: L’Assurance Contre le Chômage, by Paul Dupont. Journal of the Royal Statistical Society, 72(1):

127–128, March 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339958.

W:1909:BRBa

[611] G. H. W. Book review: Unemployment: A Problem of Industry, by W. H. Beveridge. Journal of the Royal Statistical Society, 72(1):128–131, March 1909. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339959.

B:1909:BRBf

[612] E. B. Book review: Le Problème des Retraites Ouvrières Envisagé au Point de Vue Financier, by Albert Jouve. Journal of the Royal Statistical Society, 72(1): 132–133, March 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339960.

W:1909:BRBb

[613] G. H. W. Book review: Problems of Unemployment in the London Building Trades, by Norman B. Dearle. Journal of the Royal Statistical Society, 72(1): 133–135, March 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2339961.

B:1909:BRBg

[614] A. L. B. Book review: La Grève, les Salaires et le Contrat de Travail, by D. Zolla; M. Anatole Leroy-Beaulieu. Journal of the Royal Statistical Society, 72 (1):135–136, March 1909. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339962.

M:1909:BRB

[615] H. W. M. Book review: Les Trusts et les Syndicats de Producteurs, by J. Chastin; Lycée Voltaire. Journal of the Royal Statistical Society, 72(1):136–137, March 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339963.

L:1909:BRB

[616] W. T. L. Book review: La Propriété.

Evolution de la Fortune Privée: Enquête sur la Propriété non Bâtie et Batie: Enquête Agricole, by Edmond Michel. Journal of the Royal Statistical Society, 72(1):137–138, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339964.

C:1909:BRB

[617] P. G. C. Book review: The Book of Wheat, by P. T. Dondlinger. Journal of the Royal Statistical Society, 72(1): 138–139, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339965.

K:1909:BRBa

[618] J. M. K. Book review: Die Geographische Verteilung der Getreidepreise in Indien von 1861 bis 1905, by Th. H. Engelbrecht. Journal of the Royal Statistical Society, 72(1):139–140, March 1909. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339966.

P:1909:BRBa

[619] L. L. P. Book review: A History of the Bank of England, by A. Andréadès; Christabel Meredith; H. S. Foxwell. Journal of the Royal Statistical Society, 72(1):140–142, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339967.

Y:1909:BRBb

[620] G. U. Y. Book review: Soziale Gesetzgebung und Statistik. Ein Beitrag zur Frage der Errichtung Eines Eidgenössischen Sozialstatistischen Amtes, by N. Reichesberg. Journal of the Royal Statistical Society, 72(1):142– 144, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339968.

Y:1909:BRBc

[621] G. U. Y. Book review: Essais de Statistique Morale. I. Le Suicide, by Camille Jacquart. Journal of the Royal Statistical Society, 72(1):144–145, March 1909. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339969.

P:1909:BRBb

[622] L. L. P. Book review: Outlines of the Economic History of England: A Study in Social Development, by H. O. Meredith. Journal of the Royal Statistical Society, 72(1):145–146, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339970.

W:1909:BRBc

[623] A. D. W. Book review: The Evolution of Modern Germany, by William Harbutt Dawson. Journal of the Royal Statistical Society, 72(1):146–147, March 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339971.

K:1909:BRBb

[624] J. K. Book review: School Reports and School Efficiency, by David S. Snedden; Wm. H. Allen. Journal of the Royal Statistical Society, 72(1):147–148, March 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339972.

Anonymous:1909:ONPa

[625] Anonymous. Other new publications. Journal of the Royal Statistical Society, 72(1):149–151, March 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339973.

Anonymous:1909:SEAa

[626] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 72(1): 152–160, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339974.

Anonymous:1909:QLAa

[627] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 72(1):161–186, March 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339975.

Anonymous:1909:PR

[628] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 72(1): 187–204, March 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339976.

Hutchins:1909:SWL

[629] B. L. Hutchins. Statistics of Women’s life and employment. Journal of the Royal Statistical Society, 72(2):205–247, June 1909. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340219.

Baldwin-Wiseman:1909:INC

[630] William Ralph Baldwin-Wiseman. The increase in the national consumption of water. Journal of the Royal Statistical Society, 72(2):248–303, June 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340220.

Hooker:1909:MSU

[631] R. H. Hooker. The meat supply of the United Kingdom. Journal of the Royal Statistical Society, 72(2):304–386, June 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340221.

Anonymous:1909:RCF

[632] Anonymous. Report of the Council for the financial year ended 31st December, 1908, and for the sessional year ending 15th June, 1909, presented at the seventy-fifth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, W.C., London, on the 15th of June, 1909. Journal of the Royal Statistical Society, 72(2):387–397, June 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340222.

Anonymous:1909:PSF

[633] Anonymous. Proceedings of the seventyfifth Annual General Meeting, held on 15th June, 1909. Journal of the Royal Statistical Society, 72(2):398–400, June 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340223.

C:1909:FHF

[634] P. G. C. Frederick Hendriks, F.I.A. Journal of the Royal Statistical Society, 72(2):401–403, June 1909. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340224.

W:1909:BRBd

[635] G. H. W. Book review: Earnings and Hours Inquiry. Report…, by the Board of Trade. … Earnings and Hours of Labour. … of the United Kingdom. Journal of the Royal Statistical Society, 72(2):404–408, June 1909. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340225.

B:1909:BRBh

[636] E. B. Book review: Statistical Memoranda and Charts Prepared in the Local Government Board Relating to Public Health and Social Conditions. Journal of the Royal Statistical Society, 72(2): 409–410, June 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340226.

W:1909:BRBe

[637] A. D. W. Book review: Statistik des Deutschen Reichs Band 202. Berufsund Betriebszählung vom 12 Juni 1907. Berufsstatistik. Journal of the Royal Statistical Society, 72(2):410–412, June 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340227.

W:1909:BRBf

[638] A. D. W. Book review: Austria-Hungary. Veränderungen im Stande der Gewerbe Während der Sieben Jahresperioden 1900–01 bis 1906–07. Journal of the Royal Statistical Society, 72(2):412, June 1909. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340228.

B:1909:BRBi

[639] A. L. B. Book review: Le Chômage et la Profession. Contribution à l’Étude Statistique du Chômage et de son Coefficient Professionel, by Max Lazard. Journal of the Royal Statistical Society, 72(2):412–414, June 1909. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340229.

B:1909:BRBj

[640] E. B. Book review: Unemployment:

The Results of an Investigation made in Lancashire, by S. J. Chapman; H. M. Hallsworth. Journal of the Royal Statistical Society, 72(2):414–415, June 1909. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340230.

B:1909:BRBk

[641] E. B. Book review: Insurance Against Unemployment, by David F. Schloss. Journal of the Royal Statistical Society, 72(2):415–416, June 1909. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340231.

B:1909:BRBl

[642] E. B. Book review: L’Assurance Contre le Chômage, by Maurice Bellom. Journal of the Royal Statistical Society, 72(2): 416–417, June 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340232.

J:1909:BRBa

[643] H. S. J. Book review: Le Commerce et les Commerçants, by Yves Guyot. Journal of the Royal Statistical Society, 72(2):417–420, June 1909. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340233.

J:1909:BRBb

[644] H. S. J. Book review: The Trade and Industry of Australasia, by Ben. H. Morgan. Journal of the Royal Statistical Society, 72(2):420, June 1909. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340234.

P:1909:BRBc

[645] L. L. P. Book review: Impôts Directs et Indirects sur le Revenu, by Jules Ingenbleek. Journal of the Royal Statistical Society, 72(2):420–422, June 1909. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340235.

P:1909:BRBd

[646] L. L. P. Book review: Economic Heresies, by Nathaniel Nathan. Journal of the Royal Statistical Society, 72(2): 423–425, June 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340236.

Anonymous:1909:ONPb

[647] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 72(2):425–427, June 1909. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340237.

Anonymous:1909:SEAb

[648] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 72(2): 428–438, June 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340238.

Anonymous:1909:QLAb

[649] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 72(2):439–464, June 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340239.

Paish:1909:GBC

[650] George Paish. Great Britain’s capital investments in other lands. Journal of the Royal Statistical Society, 72(3):465– 495, September 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340356.

Atkinson:1909:RPI

[651] Fred J. Atkinson. Rupee prices in India, 1870 to 1908; with an examination of the causes leading to the present high level of prices. Journal of the Royal Statistical Society, 72(3):496–573, September 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340357.

Anonymous:1909:RCCb

[652] Anonymous. Report of the census committee, 1908–09. Journal of the Royal Statistical Society, 72(3):574–593, September 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340358.

Anonymous:1909:ISI

[653] Anonymous. The International Statistical Institute at Paris. Journal of the Royal Statistical Society, 72(3):594– 603, September 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340359.

Fisher:1909:PME

[654] Irving Fisher. A practical method of estimating the velocity of circulation of money. Journal of the Royal Statistical Society, 72(3):604–618, September 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340360.

Bisset-Smith:1909:SNB

[655] G. T. Bisset-Smith. A statistical note on birth registration in Scotland previous to 1855; suggested by inquiries as to verification of birth for old age pensions. Journal of the Royal Statistical Society, 72(3):618–623, September 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340361.

Anonymous:1909:ARGa

[656] Anonymous. Agricultural returns of Great Britain, 1909. Journal of the Royal Statistical Society, 72(3):624–625, September 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340362.

A:1909:BRBa

[657] P. A. Book review: The Theory of the Construction of Tables of Mortality, by G. F. Hardy; C. Layton; E. Layton. Journal of the Royal Statistical Society, 72(3):626–627, September 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340363.

A:1909:BRBb

[658] P. A. Book review: Die Entwicklung der Reichsfinanzen, by R. van der Borght. Journal of the Royal Statistical Society, 72(3):627–629, September 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340364.

F:1909:BRBa

[659] A. W. F. Book review: Les Riches Depuis Sept Cent Ans, by G. d’Avenel. Journal of the Royal Statistical Society, 72(3):630–631, September 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340365.

P:1909:BRBe

[660] L. L. P. Book review: Salariat et Salaires, by E. Levasseur. Journal of the Royal Statistical Society, 72(3):631– 632, September 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340366.

B:1909:BRBm

[661] P. E. B. Book review: Handbuch des Wohnungswesens und der Wohnungsfrage, by Rud. Eberstadt. Journal of the Royal Statistical Society, 72(3):632– 634, September 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340367.

H:1909:BRBb

[662] E. J. H. Book review: Les Expropriations et le Prix des Terrains à Paris (1860–1900), by Maurice Halbwachs. Journal of the Royal Statistical Society, 72(3):634–636, September 1909. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340368.

H:1909:BRBc

[663] E. J. H. Book review: The Standard of Living Among Working Men’s Families in New York City, by Robert C. Chapin. Journal of the Royal Statistical Society, 72(3):637–638, September 1909. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340369.

P:1909:BRBf

[664] L. L. P. Book review: The Industrial System: An Inquiry into Earned and Unearned Income, by J. A. Hobson. Journal of the Royal Statistical Society, 72(3):638–642, September 1909. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340370.

B:1909:BRBn

[665] E. B. Book review: Royal Commission on the Poor Laws and Relief of Distress, by Constance Williams; Thomas Jones. Journal of the Royal Statistical Society, 72(3):642–643, September 1909. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340371.

B:1909:BRBo

[666] E. B. Book review: Making of our Clothes: A Case for Trade Boards, Being the Results of a Year’s Investigation into the Work of Women in London in the Tailoring, Dressmaking, and Underclothing Trades, by Carl Meyer; Clementina Black. Journal of the Royal Statistical Society, 72(3):643– 644, September 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340372.

F:1909:BRBb

[667] A. W. F. Book review: Die Bewegung des Silberpreises seit 1873, by Paul George. Journal of the Royal Statistical Society, 72(3):644–645, September 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340373.

Hesse:1909:BRB

[668] A. Hesse and A. D. W. Book review: Grundriss zum Studium der Politischen Oekonomie von Prof. Dr. J. Conrad. Vierter Teil: Statistik. Journal of the Royal Statistical Society, 72(3): 645–646, September 1909. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340374.

W:1909:BRBg

[669] A. D. W. Book review: Statistik des Deutschen Reichs. Journal of the Royal Statistical Society, 72(3):646– 648, September 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340375.

Anonymous:1909:ONPc

[670] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 72(3):648–652, September 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340376.

Anonymous:1909:SEAc

[671] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 72(3):653– 659, September 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340377.

Anonymous:1909:QLAc

[672] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 72(3):660–684, September 1909. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340378.

Anonymous:1909:VIb

[673] Anonymous. Volume information. Journal of the Royal Statistical Society, 72 (4):797–811, December 1909. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340136.

Anonymous:1909:RGP

[674] Anonymous. The recent growth of population in Western Europe; an essay in international comparison. Journal of the Royal Statistical Society, 72(4):685– 713, December 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340137.

Anonymous:1909:PN

[675] Anonymous. Proceedings on the 16th November, 1909. Journal of the Royal Statistical Society, 72(4):714– 717, December 1909. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340138.

Bowley:1909:SIC

[676] A. L. Bowley. A suggestion for the international comparison of wages by the use of the median. Journal of the Royal Statistical Society, 72(4):718–721, December 1909. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340139.

Yule:1909:AMC

[677] G. Udny Yule. The applications of the method of correlation to social and economic statistics. Journal of the Royal Statistical Society, 72(4):721– 730, December 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340140.

Anonymous:1909:IOS

[678] Anonymous. Improvement of official statistics. Journal of the Royal Statistical Society, 72(4):731–743, December 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340141.

Anonymous:1909:ARGb

[679] Anonymous. Agricultural returns of Great Britain, 1909. Journal of the Royal Statistical Society, 72(4):744– 745, December 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340142.

Anonymous:1909:SSP

[680] Anonymous. La Société de Statistique de Paris. Notes sur Paris. (À l’occasion du cinquantenaire de la Société et de la XII^{e} session de l’Institut International de Statistique). Journal of the Royal Statistical Society, 72(4):746– 748, December 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340143.

Anonymous:1909:CPP

[681] Anonymous. Census of production (1907). Preliminary tables summarising the results of the returns received under the census of Production Act, 1906. Journal of the Royal Statistical Society, 72(4):748–750, December 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340144.

J:1909:BRBc

[682] H. S. J. Book review: Earnings and Hours Inquiry. Report, by the Board of Trade into the Earnings and Hours of Labour of Workpeople of the United Kingdom. II. Clothing Trades in 1906. Journal of the Royal Statistical Society, 72(4):750–754, December 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340145.

F:1909:BRBc

[683] A. W. F. Book review: A Century of Population Growth, 1790–1900. Journal of the Royal Statistical Society, 72 (4):754–756, December 1909. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340146.

B:1909:BRBp

[684] E. B. Book review: Industrial Insurance in the United States, by Charles Richmond Henderson. Journal of the Royal Statistical Society, 72(4):756– 757, December 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340147.

B:1909:BRBq

[685] E. B. Book review: State Insurance: A Social and Industrial Need, by Frank W. Lewis. Journal of the Royal Statistical Society, 72(4):757–758, December 1909. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340148.

Anonymous:1909:ONPd

[686] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 72(4):759–763, December 1909. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340149.

Anonymous:1909:SEAd

[687] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 72(4):764– 771, December 1909. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340150.

Anonymous:1909:QLAd

[688] Anonymous. Quarterly list of additions to the library. Journal of the Royal Statistical Society, 72(4):772–796, December 1909. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340151.

Anonymous:1910:ESU

[689] Anonymous. Errata: On the smaller urban districts of England and Wales. Journal of the Royal Statistical Society, 74(1):79, December 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339984. See [187].